### v1

# Some Constructions on Quantum Principal Bundles

Abstract.
This paper aims to prove four statements in the theory of quantum principal bundles:
  1. (1)
    The universal differential envelope $\ast$–calculus of a matrix (compact) Lie group, for the classical bicovariant $\ast$–First Order Differential Calculus, is the algebra of differential forms.
  2. (2)
    An example of a quantum principal bundle in which the space of base forms is not generated by the base space.
  3. (3)
    The group isomorphism between convolution-invertible maps and covariant left module isomorphisms at the level of differential calculus
  4. (4)
    The way the maps $\{T^{V}_{k}\}$ from Remark 3.1 look in differential geometry.
Furthermore, we are going to present an example of the theory of the reference [1] in a special classical/quantum hybrid principal bundle.
1. Universal Differential Envelope $\ast$–calculus in Differential Geometry
A compact matrix quantum group (or a just a quantum group) will be denoted by $\mathcal{G}$, and its dense $\ast$–Hopf (sub)algebra will be denoted by
(1) $H^{\infty}:=(H,\cdot,\mathbbm{1},\Delta,\epsilon,S,\ast),$
where $\Delta$ is the coproduct, $\epsilon$ is the counity and $S$ is the coinverse. The space $H^{\infty}$ shall be treated as the algebra of all polynomial functions defined on $\mathcal{G}$. In the same way a (smooth right) $\mathcal{G}$–corepresentation on a $\mathbb{C}$–vector space $V$ is a linear map
$\delta^{V}:V\longrightarrow V\otimes H$
such that
(2) $(\mathrm{id}_{V}\otimes\epsilon)\circ\delta^{V}\cong\mathrm{id}_{V}$
and
(3) $(\mathrm{id}_{V}\otimes\Delta)\circ\delta^{V}=(\delta^{V}\otimes\mathrm{id}_{H})\circ\delta^{V}.$
We say that the corepresentation is finite–dimensional if $\dim_{\mathbb{C}}(V)<|\mathbb{N}|$. $\delta^{V}$ usually receives the name of the (right) coaction of $\mathcal{G}$ on $V$. It is worth mentioning that in the general theory ([2]), equation (2) is not necessary.
The set of all corepresentation morphisms between two corepresentations $\delta^{V}$, $\delta^{W}$ will be denoted as
(4) $\textsc{Mor}(\delta^{V},\delta^{W}),$
A $\mathcal{G}$–corepresentation $\delta^{V}$ is reducible if there exists a non–trivial subspace $L$ $(L\not=\{0\},V)$ such that $\delta^{V}(L)=L\otimes H$ and $\delta^{V}$ is unitary if viewed as an element of $B(V)\otimes H$ (with $B(V):=\{f:V\longrightarrow V\mid f\mbox{ is linear}\}$) is unitary. Of course, for the last definition, it is necessary an inner product $\langle$-|$\rangle$ - | - ⟩ on $V$. In [2], Woronowicz proved that for every finite–dimensional $\mathcal{G}$–corepresentation on $V$ there is an inner product $\langle$-|$\rangle$ - | - ⟩ (not necessarily unique) making the corepresentation unitary. Thus, henceforth we are going to consider that every finite–dimensional $\mathcal{G}$–corepresentation is unitary. In [2] one can find a proof of the following theorem.
Theorem 1.1.
Let $\mathcal{T}$ be a complete set of mutually non–equivalent irreducible unitary (necessarily finite–dimensional) $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$ (the trivial corepresentation on $\mathbb{C}$). For any $\delta^{V}$in $\mathcal{T}$ that coacts on $(V,\langle-|-\rangle)$,
(5) $\delta^{V}(e_{i})=\sum^{n_{V}}_{j=1}e_{j}\otimes g^{V}_{ji},$
where $\{e_{i}\}^{n_{V}}_{i=1}$ is an ortonormal basis of $V$ and $\{g^{V}_{ij}\}^{n_{V}}_{i,j=1}$subseteq $H$. Then $\{g^{V}_{ij}\}_{\delta^{V},i,j}$ is a linear basis of $H$, where the index $\delta^{V}$ runs on $\mathcal{T}$ and $i$, $j$ run from $1$ to $n_{V}=\mathrm{dim}_{\mathbb{C}}(V)$.
For every $\delta^{V}$in $\mathcal{T}$, the set $\{g^{V}_{ij}\}^{n_{V}}_{ij=1}$ satisfies
(6) $\phi(g^{\alpha}_{ij})=\sum^{n_{\alpha}}_{k=1}g^{\alpha}_{ik}\otimes g^{\alpha}_{kj},\quad\kappa(g^{\alpha}_{ij})=g^{\alpha\,\ast}_{ji},\quad\sum^{n_{\alpha}}_{k=1}g^{\alpha}_{ik}g^{\alpha\,\ast}_{jk}=\sum^{n_{\alpha}}_{k=1}g^{\alpha\,\ast}_{ki}g^{\alpha}_{kj}=\delta_{ij}\mathbbm{1},\quad\epsilon(g^{\alpha}_{ij})=\delta_{ij}$
with $\delta_{ij}$ being the Kronecker delta, among other properties [2].
Let $(\Gamma,d)$ be a bicovariant $\ast$–First Order Differential Calculus ($\ast$–FODC) over $\mathcal{G}$ ([3, 6]) and consider the $\mathbb{C}$–vector space given by
(7) $\mathfrak{qg}^{\#}:=\{\theta\in\Gamma\mid\Phi_{\Gamma}(\theta)=\mathbbm{1}\otimes\theta\}=\mathbbm{1}\otimes{\mathrm{Ker}(\epsilon)\over\mathcal{R}}\cong{\mathrm{Ker}(\epsilon)\over\mathcal{R}}.$
This space allows to consider the quantum germs map
(8) $\pi:H$ $\longrightarrow\mathfrak{qg}^{\#}$
$g$ $\longmapsto[g-\epsilon(g)\mathbbm{1}]_{\mathcal{R}},$
where $[h]_{\mathcal{R}}$ denotes the equivalence class of the element $h$in $H$. The map $\pi$ has several useful properties, for example the restriction map $\pi|_{\mathrm{Ker}(\epsilon)}$ is surjective and
(9) $\ker(\pi)=\mathcal{R}\oplus\mathbb{C}\mathbbm{1},$ $\qquad\;\;\,\pi(g)=S(g^{1})dg^{2},$ $\quad dg=g^{(1)}\pi(g^{(2)}),$
$\pi(g)=-(dS(g^{(1)}))g^{(2)},$ $\quad dS(g)=-\pi(g^{(1)})S(g^{(2)}),$ $\quad\pi(g)^{\ast}=-\pi(S(g)^{\ast})$
for all $g$in $H$ [6]. It is worth mentioning that the canonical right action ${{}_{\Gamma}}\Phi$ of the $\ast$–FODC leaves $\mathfrak{qg}^{\#}$ invariant ([6]); so
(10) $\mathrm{ad}:={{}_{\Gamma}}\Phi|_{\mathfrak{qg}^{\#}}:\mathfrak{qg}^{\#}\longrightarrow\mathfrak{qg}^{\#}\otimes H$
is a $\mathcal{G}$–corepresentation and it fulfills [6]
(11) $\mathrm{ad}\circ\pi=(\pi\otimes\mathrm{id}_{H})\circ\mathrm{Ad}.$
On the other hand there is a right $H$–module structure on $\mathfrak{qg}^{\#}$ given by
(12) $\theta\circ g=S(g^{(1)})\theta g^{(2)}=\pi(hg-\epsilon(h)g)$
for every $\theta=\pi(h)$in $\mathfrak{qg}^{\#}$; and it satisfies [6]
$(\theta\circ g)^{\ast}=\theta^{\ast}\circ S(g)^{\ast}.$
Let $(\Gamma,d)$ be a $\ast$–FODC over $\mathcal{G}$. Consider the space
(13) $\Gamma^{\wedge}:=\otimes^{\bullet}_{H}\Gamma/\mathcal{Q}\,,\quad\otimes^{\bullet}_{H}\Gamma:=\bigoplus_{k}(\otimes^{k}_{H}\Gamma)\quad\mbox{ with }\quad\otimes^{k}_{H}\Gamma:=\underbrace{\Gamma\otimes_{H}\cdots\otimes_{H}\Gamma}_{k\;times},$
where $\mathcal{Q}$ is the bilateral ideal of $\otimes^{\bullet}_{H}\Gamma$ generated by
(14) $\sum_{i}dg_{i}\otimes_{H}dh_{i}\quad\mbox{ such that }\quad\sum_{i}g_{i}\,dh_{i}=0,$
for $g_{i}$, $h_{i}$in $H$. On the other hand, for a given $t=\vartheta_{1}\cdots\vartheta_{n}$in $\Gamma^{\wedge\,n}$ the linear map
(15) $d:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}$
[…]
Let $(\Gamma,d)$ be a bicovariant $\ast$–FODC. Then the coproduct can be extended to a graded differential $\ast$–algebra morphism
(16) $\Delta:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}\otimes\Gamma^{\wedge}.$
The counit and the coinverse can also be extended to $(\Gamma^{\wedge},d,\ast)$. These maps define a graded differential $\ast$–Hopf algebra structure
(17) $\Gamma^{\wedge\,\infty}:=(\Gamma^{\wedge},\cdot,\mathbbm{1},\Delta,\epsilon,S,d,\ast),$
on $\Gamma^{\wedge}$ which extends $H^{\infty}=(H,\cdot,\mathbbm{1},\Delta,\epsilon,S,d,\ast)$ [4]. In particular,
(18) $\epsilon:\Gamma^{\wedge}\longrightarrow\mathbb{C}$
given by $\epsilon|_{H}:=\epsilon$, $\epsilon|_{\Gamma^{\wedge k}}:=0$ for $k\geq 1$.
Now one can consider the right adjoint action of $\Gamma^{\wedge}$ by
(19) $\mathrm{Ad}:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}\otimes\Gamma^{\wedge}$
such that $\mathrm{Ad}(\vartheta)=(-1)^{\partial\vartheta^{(1)}\partial\vartheta^{(2)}}\vartheta^{(2)}\otimes S(\vartheta^{(1)})\vartheta^{(3)},$ where $\partial x$ denotes the grade of $x$ and $(\mathrm{id}_{\Gamma^{\wedge}}\otimes\Delta)\Delta(\vartheta)=(\Delta\otimes\mathrm{id}_{\Gamma^{\wedge}})\Delta(\vartheta)=\vartheta^{(1)}\otimes\vartheta^{(2)}\otimes\vartheta^{(3)}.$ Clearly, $\mathrm{Ad}$ extends the right adjoint coaction $\mathrm{Ad}$ of $H$.
Let us define
(20) $\mathfrak{qg}^{\#\wedge}=\otimes^{\bullet}\mathfrak{qg}^{\#}/A^{\wedge},\quad\otimes^{\bullet}\mathfrak{qg}^{\#}:=\oplus_{k}(\otimes^{k}\mathfrak{qg}^{\#})\;\mbox{ with }\;\otimes^{k}\mathfrak{qg}^{\#}:=\underbrace{\mathfrak{qg}^{\#}\otimes\cdots\otimes\mathfrak{qg}^{\#}}_{k\;times},$
where $A^{\wedge}$ is the bilateral ideal of $\otimes^{\bullet}\mathfrak{qg}^{\#}$ generated by
$\pi(g^{(1)})\otimes\pi(g^{(2)})$
with $g$in $\mathcal{R}$ [4]. The space $\mathfrak{qg}^{\#\wedge}$ is a graded differential $\ast$–subalgebra ([4]) and it satisfies $\mathfrak{qg}^{\#\wedge}=\{\theta\in\Gamma^{\wedge}\mid\Phi_{\Gamma^{\wedge}}(\theta)=\mathbbm{1}\otimes\theta\}$. Also one can extend the right $H$–module structure of $\mathfrak{qg}^{\#}$ (see equation (12)) to $\mathfrak{qg}^{\#^{\wedge}}$ by means of
(21) $1\circ g=\epsilon(g),\quad(\theta_{1}\theta_{2})\circ g=(\vartheta_{1}\circ g^{(1)})(\theta_{2}\circ g^{(2)}).$
It is worth mentioning that
(22) $d\pi(g)=-\pi(g^{(1)})\pi(g^{(2)}).$
According to [4] the following identification holds:
(23) $\Gamma^{\wedge}=H\otimes\mathfrak{qg}^{\#\wedge}.$
The following example is one of the purposes of this paper
Example 1.2.
Let $G$ be a compact matrix Lie group ($G\subseteq M_{k}(\mathbb{C}))$), and let $\mathcal{G}$ be its associated quantum group [2]. If $H^{\infty}=(H,\cdot,\mathbbm{1},\Delta,\epsilon,S,\ast,)$ is its dense $\ast$–Hopf algebra, then $H$ is the $\ast$–algebra generated by the smooth $\mathbb{C}$–valued functions
$w_{ij}:G$ $\longrightarrow\mathbb{C}$
$A$ $\longmapsto a_{ij}.$
with $A=(a_{ij})$. The coproduct of an element $g$in $H$ is defined by
$\Delta(g):G\times G\longmapsto\mathbb{C},\qquad\Delta(g)(A,B)=g(AB)\quad\mbox{with }\quad A,\,B\,\in\,G.$
In particular, we have $\Delta(w_{ij})=\displaystyle\sum^{n}_{k=1}w_{ik}\otimes w_{kj}.$ The counit and the coinverse are defined as follows:
$\epsilon:H\longrightarrow\mathbb{C},\qquad g\longmapsto g(e),$
where $e$in $G$ is the identity element; and
$S(g):G\longrightarrow\mathbb{C},\qquad A\longmapsto S(g)(A)=g(A^{-1})$
for all $g$in $H$.
Since $G$ is parallelizable, we have the identification $G\times\mathfrak{g}=TG$, where $\mathfrak{g}$ is the Lie algebra of $G$. This diffeomorphism can be given by
(24) $(A,v)\longmapsto(dL_{A})_{e}(v),$
where $(dL_{A})_{e}$ is the differential at $e$ of the left product on $G$ with $A$; and by
(25) $(A,v)\longmapsto(dR_{A})_{e}(v),$
where $(dR_{A})_{e}$ is the differential at $e$ of the right product on $G$ with $A$.
Under equation (24) we have $dg=g^{(1)}\otimes dg^{(2)}_{e}$ for all $g$in $H$, with $dh_{e}$ the differential of $h$in $H$ at $e$. By defining
$\mathfrak{h}:=\mathrm{span}_{\mathbb{C}}\{dg_{e}\mid g\,\in\,H\}$
we obtain that
$(\Gamma=H\otimes\mathfrak{h},d)$
is a $\ast$–FODC over $\mathcal{G}$, where $d$ is the usual differential of smooth $\mathbb{C}$–valued functions. Moreover, it is left–covariant under the map $\Phi_{\Gamma}=\Delta\otimes\mathrm{id}_{\mathfrak{h}}$.
On the other hand, under equation (25), we have $dg=dg^{(1)}_{e}\otimes g^{(2)}$ for all $g$in $H$. Thus
$(\Gamma,d)=(\mathfrak{h}\otimes H,d)$
and it follows that $(\Gamma,d)$ is right–covariant by the map ${{}_{\Gamma}}\Phi=\mathrm{id}_{\mathfrak{h}}\otimes\Delta.$ Hence, $(\Gamma,d)$ is bicovariant. According to [4, 6], there exists a right $H$–ideal $\mathcal{R}\subseteq\mathrm{Ker}(\epsilon)$ such that
$\{\theta\in\Gamma\mid\Phi_{\Gamma}(\theta)=\mathbbm{1}\otimes\theta\}=\mathbbm{1}\otimes{\mathrm{Ker}(\epsilon)\over\mathcal{R}}={\mathrm{Ker}(\epsilon)\over\mathcal{R}}.$
However, since $\mathfrak{h}\cong\mathbbm{1}\otimes\mathfrak{h}=\{\theta\in\Gamma\mid\Phi_{\Gamma}(\theta)=\mathbbm{1}\otimes\theta\}$ it follows that
$\mathfrak{h}={\mathrm{Ker}(\epsilon)\over\mathcal{R}}=\mathfrak{qg}^{\#}.$
We claim that $\mathcal{R}=\mathrm{Ker}^{2}(\epsilon):=\{\displaystyle\sum^{n}_{i=1}a_{i}\,b_{i}\mid a_{i},\,b_{i}\,\in\,\mathrm{Ker}(\epsilon)\,\mbox{ for some }\,n\,\in\,\mathbb{N}\}$. In fact, consider the quantum germs map
$\pi:\mathrm{Ker}(\epsilon)$ $\longrightarrow{\mathrm{Ker}(\epsilon)\over\mathcal{R}}$
$g$ $\longmapsto[g]_{\mathcal{R}}.$
Then, by equation (9), we obtain $\pi(g)=S(g^{(1)})dg^{(2)}=dg_{e}$. If $\mathfrak{g}^{\#}_{\mathbb{C}}$ is the complexification of the dual space of $\mathfrak{g}$, it is well–known that for $X=\{f\in C^{\infty}_{\mathbb{C}}(G)\mid\epsilon(f)=0\}$ we have
$\mathfrak{g}^{\#}_{\mathbb{C}}={X\over X^{2}}$
with $X^{2}=\{\displaystyle\sum^{n}_{i=1}a_{i}\,b_{i}\mid a_{i},\,b_{i}\,\in\,X\,\mbox{ for some }\,n\,\in\,\mathbb{N}\}$. Moreover, for every $f$in $X$,
$df_{e}=[f]_{X^{2}},$
where $[f]_{X^{2}}$ denotes the equivalence class of $f$ in $X/X^{2}$. In this way, for every $g$in $\mathrm{Ker}(\epsilon)$ we get $dg_{e}=[g]_{\mathcal{R}}=[g]_{X^{2}}$. By this relation, $g$in $\mathcal{R}$ if and only if $g$in $=X^{2}\cap\mathrm{Ker}(\epsilon)=\mathrm{Ker}^{2}(\epsilon),$ which proves that $\mathcal{R}=\mathrm{Ker}^{2}(\epsilon)$.
By definition, $\mathfrak{h}\subseteq\mathfrak{g}_{\mathbb{C}}$. Since $\mathfrak{g}_{\mathbb{C}}\subseteq M_{k}(\mathbb{C})$, there exists a linear basis $\{v_{i}\}$ of $\mathfrak{g}_{\mathbb{C}}$ composed of linear combinations of the canonical linear basis $\{E_{ij}\}$ of $M_{k}(\mathbb{C})$ ($E_{ij}$ is the matrix with $1$ in the $(i,j)$ position and 00 elsewhere). Then $(dw_{ij})_{e}(v_{k})$ corresponds to the $(i,j)$ position of the matrix $v_{k}$. Consequently, there exist elements $\{(dg_{i})_{e}\}$, which are linear combinations of $(dw_{ij})_{e}$in $\mathfrak{h}$, such that $(dg_{i})_{e}(v_{j})=\delta_{ij}$ with $\delta_{ij}$ being the Kronecker delta. Thus, $\mathfrak{g}^{\#}_{\mathbb{C}}=\mathfrak{h}=\mathfrak{qg}^{\#}$ and we obtain
$(\Gamma,d)=(H\otimes\mathfrak{g}^{\#}_{\mathbb{C}},d).$
This implies that $(\Gamma,d)$ is a $\ast$–subFODC of the $\ast$–FODC over $C^{\infty}_{\mathbb{C}}(G)$
$(C^{\infty}_{\mathbb{C}}(G)\otimes\mathfrak{g}^{\#}_{\mathbb{C}},d)$
of $\mathbb{C}$–valued differential $1$–forms of $G$. Notice that it suffices to take convergent sequences $\{g_{i}\}^{\infty}_{i=1}$subseteq $H$ in $(\Gamma,d)$ to recover $(C^{\infty}_{\mathbb{C}}(G)\otimes\mathfrak{g}^{\#}_{\mathbb{C}},d)$, although this lies beyond our purely geometric–algebraic approach. It is worth mentioning that for the specific form $\Gamma=H\otimes\mathfrak{g}^{\#}_{\mathbb{C}}$, the right covariance map ${{}_{\Gamma}}\Phi$ is given by [6]
${{}_{\Gamma}}\Phi(g\otimes\pi(h))=g^{(1)}\otimes\pi(h^{(2)})\otimes g^{(2)}S(h^{(1)})h^{(1)}$
and hence the $\mathcal{G}$–corepresentation $\mathrm{ad}$ (see equation (10)) coincides with the dualization of the right adjoint action of $G$ on $\mathfrak{g}_{\mathbb{C}}$.
By the Leibniz rule, we have
$\pi(hg)=\pi(h)\epsilon(g)+\epsilon(h)\pi(g)$
and the right $H$–module structure on $\mathfrak{g}^{\#}_{\mathbb{C}}$, given by (see equation (12)) $\pi(h)\circ g=\pi(hg-\epsilon(h)g)$ simplifies to
$\pi(h)\circ g=\epsilon(g)\pi(h).$
Now, consider the universal differential envelope $\ast$–calculus
$(\Gamma^{\wedge},d,\ast)$
of $(\Gamma,d)$. If $g=a\,b$in $\mathrm{Ker}^{2}(\epsilon)$, then by equation (12), we obtain
$\pi(g^{(1)})\otimes\pi(g^{(2)})=\pi(a^{(1)}b^{(1)})\otimes\pi(a^{(2)}b^{(2)})$ $=$ $\pi(a^{(1)})\circ b^{(1)}\otimes\pi(a^{(2)})\circ b^{(2)}$
$+$ $\epsilon(a^{(1)})\pi(b^{(1)})\otimes\pi(a^{(2)})\circ b^{(2)}$
$+$ $\pi(a^{(1)})\circ b^{(1)}\otimes\epsilon(a^{(2)})\pi(b^{(2)})$
$+$ $\epsilon(a^{(1)})\pi(b^{(1)})\otimes\epsilon(a^{(2)})\pi(b^{(2)})$
$=$ $\pi(b)\otimes\pi(a)+\pi(a)\otimes\pi(b).$
and therefore (see equation (20))
$\mathfrak{g}^{\#\,\wedge}_{\mathbb{C}}=\otimes^{\bullet}\mathfrak{g}^{\#}_{\mathbb{C}}/A^{\wedge}=\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}}.$
Thus (see equation (23))
$(\Gamma^{\wedge}=H\otimes\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}},d,\ast).$
By the axioms of the exterior derivative, we can conclude that $d:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}$ is precisely the exterior derivative. Hence $(\Gamma^{\wedge},d,\ast)$ is a graded differential $\ast$–subalgebra of the graded differential $\ast$–algebra
$(C^{\infty}_{\mathbb{C}}(G)\otimes\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}},d,\ast)$
of $\mathbb{C}$–valued differential forms of $G$. As before, it suffices to consider convergent sequences $\{g_{i}\}^{\infty}_{i=1}$subseteq $H$ in $(\Gamma^{\wedge},d,\ast)$ to obtain $(C^{\infty}_{\mathbb{C}}(G)\otimes\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}},d,\ast),$ although this lies beyond our purely geometric–algebraic approach.
The map $\Delta$ is a graded differential $\ast$–morphism. Thus, we have
$\Delta(g_{0}\,\wedge dg_{1}\wedge\cdots\wedge dg_{n})=\Delta(g_{0})\cdot d\Delta(g_{1})\cdots d\Delta(g_{n}).$
Moreover,
$\epsilon(g_{0}\,\wedge dg_{1}\wedge\cdots\wedge dg_{n})=\epsilon(g_{0})d\epsilon(g_{1})\cdots d\epsilon(g_{n})=0,$
as expected. Since $S$ is a graded antimultiplicative map that commutes with $d$, it follows that
$S(g_{0}\,\wedge dg_{1}\wedge\cdots\wedge dg_{n})=(-1)^{n(n-1)\over 2}dS(g_{n})\wedge\cdots\wedge dS(g_{1})\cdot S(g_{0}).$
2. An example about the space of base forms.
Let $(B,\cdot,\mathbbm{1},\ast)$ be a quantum space and let $\mathcal{G}$ be a quantum group. A quantum principal $\mathcal{G}$–bundle over $B$ (abbreviated ”qpb”) is a quantum structure formally represented by the triplet
(26) $\zeta=(P,B,\Delta_{P}),$
where $(P,\cdot,\mathbbm{1},\ast)$ is called the quantum total space, and $(B,\cdot,\mathbbm{1},\ast)$ a quantum subspace, which receives the name quantum base space. Furthermore,
$\Delta_{P}:P\longrightarrow P\otimes H$
is a $\ast$–algebra morphism that satisfies
  1. (1)
    $\Delta_{P}$ is a $\mathcal{G}$–corepresentation.
  2. (2)
    $\Delta_{P}(x)=x\otimes\mathbbm{1}$ if and only if $x$in $B$.
  3. (3)
    The linear map $\beta:P\otimes P\longrightarrow P\otimes H$ given by
    $\beta(x\otimes y):=x\cdot\Delta_{P}(y):=(x\otimes\mathbbm{1})\cdot\Delta_{P}(y)$
    is surjective.
In general, one does not need a quantum group, a $\ast$–Hopf algebra suffices [5, 6]. Given $\zeta$ a qpb over $B$, a differential calculus on it is:
  1. (1)
    A graded differential $\ast$–algebra $(\Omega^{\bullet}(P),d,\ast)$ generated by $\Omega^{0}(P)=P$ (quantum differential forms of $P$).
  2. (2)
    A bicovariant $\ast$–FODC $(\Gamma,d)$ over $\mathcal{G}$.
  3. (3)
    The map $\Delta_{P}$ is extendible to a graded differential $\ast$–algebra morphism
    $\Delta_{\Omega^{\bullet}(P)}:\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}.$
    Here we have considered that $\otimes$ is the tensor product of graded differential $\ast$–algebras.
Note that if $\Delta_{\Omega^{\bullet}(P)}$ exists, it is unique because our graded differential $\ast$–algebras are generated by its degree 00 elements. Furthermore, $\Delta_{\Omega^{\bullet}(P)}$ is a graded differential $\Gamma^{\wedge}$–corepresentation on $\Omega^{\bullet}(P)$ [5]. In this setting, the space of horizontal forms is defined as
(27) $\mathrm{Hor}^{\bullet}P\,:=\{\varphi\in\Omega^{\bullet}(P)\mid\Delta_{\Omega^{\bullet}(P)}(\varphi)\,\in\,\Omega^{\bullet}(P)\otimes H\},$
and it is a graded $\ast$–subalgebra of $\Omega^{\bullet}(P)$ [6]. Since $\Delta_{\Omega^{\bullet}(P)}(\mathrm{Hor}^{\bullet}P)\subseteq\mathrm{Hor}^{\bullet}P\otimes H,$ the map
(28) $\Delta_{\mathrm{Hor}}:=\Delta_{\Omega^{\bullet}(P)}|_{\mathrm{Hor}^{\bullet}P}:\mathrm{Hor}^{\bullet}P\longrightarrow\mathrm{Hor}^{\bullet}P\otimes H$
is $\mathcal{G}$–corepresentation on $\mathrm{Hor}^{\bullet}P$. Also, one can define the space of base forms (quantum differential forms of $B$) as
(29) $\Omega^{\bullet}(B):=\{\mu\in\Omega^{\bullet}(P)\mid\Delta_{\Omega^{\bullet}(P)}(\mu)=\mu\otimes\mathbbm{1}\}.$
The space of base forms is a graded differential $\ast$–subalgebra of $(\Omega^{\bullet}(P),d,\ast)$, and in general, it is not generated by $B$. In fact, the following example, one of the main points of this paper, illustrates a situation in which $(\Omega^{\bullet}(B),d,\ast)$ is not generated by $B$.
Example 2.1.
Let $\mathcal{G}$ be the quantum group associated with $U(1)$ and let
$(H=\mathbb{C}[z,z^{\ast}],\cdot,\mathbbm{1},\Delta,\epsilon,S,\ast)$
[…]
$\zeta=(P:=H,B,\Delta_{P}:=\Delta),$
which is a quantum principal $\mathcal{G}$–bundle over $B$.
For quantum differential forms of $P$, we will use the universal differential envelope $\ast$–calculus $(\Omega^{\bullet}(P),d,\ast)$ presented in Example 1.2 for $U(1)$. Recall that
${{}_{\mathrm{inv}}}\Omega^{1}(P):={\mathrm{Ker}(\epsilon)\over\mathrm{Ker}^{2}(\epsilon)}=\mathfrak{u}(1)^{\#}_{\mathbb{C}}$
is the complexification of the dual space of $\mathfrak{u}(1)$, the Lie algebra of $U(1)$. Of course in this case, $\Omega^{\bullet}(P)$ has no elements of degree $n\geq 2$, and
${{}_{\mathrm{inv}}}\Omega^{1}(P)=\mathrm{span}_{\mathbb{C}}\{\pi(z)=dz_{e}\}.$
In accordance with [6], the set $\{\pi(z)\}$ is a left $P$–basis of $\Omega^{1}(P)$.
For quantum differential forms of $\mathcal{G}$, we will take the universal differential envelope $\ast$–calculus of the bicovariant $\ast$–FODC over $\mathcal{G}$ defined by
$\mathcal{R}=\mathrm{Ker}(\epsilon).$
[…]
Finally, we define
$\Delta_{\Omega^{\bullet}(P)}:\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)\otimes H$
to coincide with $\Delta_{P}$ on the degree 00 case, and for degree $1$ elements,
$\Delta_{\Omega^{\bullet}(P)}(g\,\pi(z))=\Delta(g)(\pi(z)\otimes\mathbbm{1})=g^{(1)}\,\pi(z)\otimes g^{(2)}$
for all $g$in $P$. These constructions provide a differential calculus on $\zeta$. Consequently,
$\Omega^{\bullet}(B)=B\oplus\Omega^{1}(B)\qquad\mbox{ with }\qquad\Omega^{1}(B)=\{\lambda\,\pi(z)\mid\lambda\in\mathbb{C}\}.$
However, this space is not generated by $B$ because $dB=\{0\}$.
3. The Group Isomorphism
In this section, we will consider
Remark 3.1.
Let $\zeta$ be a qpb. Henceforth, we are going to consider that the quantum base space $(B,\cdot,\mathbbm{1},\ast)$ is a $C^{\ast}$–algebra (or it can be completed to a $C^{\ast}$–algebra). Under this assumption and in accordance with [7], for every $\delta^{V}$in $\mathcal{T}$ there exists
$\{T^{\mathrm{L}}_{k}\}^{d_{V}}_{k=1}\subseteq\textsc{Mor}(\delta^{V},\Delta_{P})$
for some $d_{V}$in $\mathbb{N}$ such that
(30) $\sum^{d_{V}}_{k=1}x^{V\,\ast}_{ki}x^{V}_{kj}=\delta_{ij}\mathbbm{1},$
with $x^{V}_{ki}:=T^{\mathrm{L}}_{k}(e_{i})$, where $\mathcal{T}$ is a complete set of mutually non–equivalent irreducible (necessarily finite–dimensional) $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$ (the trivial corepresentation on $\mathbb{C}$) and $\{e_{i}\}^{n_{V}}_{i=1}$ is the orthonormal basis of $V$ shown in Theorem 1.1.
Let $\zeta=(P,B,\Delta_{P})$ be a qpb. The surjective map $\beta$ can be used to define the linear isomorphism
$\widetilde{\beta}:P\otimes_{B}P\longrightarrow P\otimes H$
such that $\widetilde{\beta}(x\otimes_{B}y)=\beta(x\otimes y)=(x\otimes\mathbbm{1})\cdot\Delta_{P}(y).$ [6]. The degree zero quantum translation map is defined as
(31) $\mathrm{qtrs}:H\longrightarrow P\otimes_{B}P,$
such that $\mathrm{qtrs}(g)=\widetilde{\beta}^{-1}(\mathbbm{1}\otimes g).$ Explicitly, by taking the linear basis $\{g^{V}_{ij}\}_{\delta^{V},i,j}$ (see Theorem 1.1 and Remark 3.1), we have
(32) $\mathrm{qtrs}(g^{V}_{ij})=\sum^{d_{V}}_{k=1}x^{V\,\ast}_{ki}\otimes_{B}x^{V}_{kj}.$
In particular, since $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$ we have that $\mathbbm{1}$in $\{g^{V}_{ij}\}_{\delta^{V},i,j}$ and
$\mathrm{qtrs}(\mathbbm{1})=\mathbbm{1}\otimes_{B}\mathbbm{1}.$
We can extend $\mathrm{qtrs}$ to
(33) $\widetilde{\mathrm{qtrs}}:P\otimes H\longrightarrow P\otimes_{B}P$
by means of $\widetilde{\mathrm{qtrs}}(x\otimes g^{V}_{ij})=x\,\mathrm{qtrs}(g^{V}_{ij})=\displaystyle\sum^{d_{\alpha}}_{k=1}x\,x^{V\,\ast}_{ki}\otimes_{B}x^{V}_{kj}.$ A direct calculation shows that $\widetilde{\beta}$ and $\widetilde{\mathrm{qtrs}}$ are mutually inverse.
Throughout the various computations of this paper, we shall use the symbolic notation
(34) $\mathrm{qtrs}(g)=[g]_{1}\otimes_{B}[g]_{2}.$
Now we shall assume that $\zeta=(P,B,\Delta)$ is endowed with a differential calculus. In this situation $\widetilde{\beta}$ has a natural extension to
(35) $\widetilde{\beta}:\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}$
given by $\widetilde{\beta}(w_{1}\otimes_{\Omega^{\bullet}(B)}w_{2})=(w_{1}\otimes\mathbbm{1})\cdot\Delta_{\Omega^{\bullet}(P)}(w_{2}),$ where the tensor product on the image is the tensor product of graded differential $\ast$–algebras. According to [9], this map is bijective.
On the other hand, taking a real qpc $\omega$ (which always exists [5]) and in accordance with [9], we can extend $\mathrm{qtrs}$ to
(36) $\mathrm{qtrs}:\Gamma\longrightarrow\left(\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P)\right)^{1}$
by means of
(37) $\mathrm{qtrs}(\theta)=\mathbbm{1}\otimes_{\Omega^{\bullet}(B)}\omega(\theta)-(m_{\Omega}\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{P})(\omega\otimes\mathrm{qtrs})\mathrm{ad}(\theta)$
when $\theta$in $\mathfrak{qg}^{\#}$, where
$m_{\Omega}:\Omega^{\bullet}(P)\otimes\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)$
is the product map; and
(38) $\mathrm{qtrs}(\vartheta\upsilon):=(-1)^{\partial\vartheta\,\partial[\upsilon]_{1}}[\upsilon]_{1}\,\mathrm{qtrs}(\vartheta)\,[\upsilon]_{2}$
for $\vartheta$in $H$, $\upsilon$in $\mathfrak{qg}^{\#}$ or $\vartheta$in $\mathfrak{qg}^{\#}$, $\upsilon$in $H$, with $\mathrm{qtrs}(\upsilon)=[\upsilon]_{1}\otimes_{\Omega^{\bullet}(B)}[\upsilon]_{2}$, where $\partial\vartheta$ is the grade of $\vartheta$. A direct calculation shows that the corresponding extended map $\widetilde{\mathrm{qtrs}}$ is the right inverse of $\widetilde{\beta}$ for degree $1$ and since $\widetilde{\beta}$ is bijective, we get that $\widetilde{\mathrm{qtrs}}$ is actually its inverse. It is worth mentioning that even when apparently the definition of $\mathrm{qtrs}$ depends on the real qpc $\omega$ chosen, by the uniqueness of the inverse function, the last result tells us that $\mathrm{qtrs}$ is independent of this choice. Since $\widetilde{\beta}$ commutes with the corresponding differential maps, it follows that ([9])
Proposition 3.2.
We have
$\mathrm{qtrs}\circ d=d_{\otimes^{\bullet}}\circ\mathrm{qtrs},$
where $d_{\otimes^{\bullet}}$ is the differential map of $\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P)$.
Now let us take the universal graded differential $\ast$ –algebra $(\otimes^{\bullet}_{H}\Gamma,d_{\otimes_{H}},\ast)$ of $(\Gamma,d)$ [6]. The quantum translation map can be extended naturally to $\otimes^{\bullet}_{H}\Gamma$ by means of
(39) $\mathrm{qtrs}(\vartheta\otimes_{H}\upsilon):=(-1)^{\partial\vartheta\,\partial[\upsilon]_{1}}\,[\upsilon]_{1}\,\mathrm{qtrs}(\vartheta)\,[\upsilon]_{2}$
if $\mathrm{qtrs}(\upsilon)=[\upsilon]_{1}\otimes_{\Omega^{\bullet}(B)}[\upsilon]_{2}$. By induction and Proposition 3.2 it can be proved that
(40) $\mathrm{qtrs}(dg_{0}\otimes_{H}dg_{1}\otimes_{H}...\otimes_{H}dg_{k})=d_{\otimes^{\bullet}}(\mathrm{qtrs}(g_{0}\,dg_{1}\otimes_{H}...\otimes_{H}dg_{k}))$
for $g_{1}$,…, $g_{k}$in $H$. Now by considering the bilateral ideal $\mathcal{Q}$ generated by the relation presented on equation (14), a direct calculation shows that $\mathrm{qtrs}(\mathcal{Q})=0$ and therefore, $\mathrm{qtrs}$ can be defined on the universal differential envelope $\ast$–calculus
(41) $\mathrm{qtrs}:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P).$
As before, the corresponding extended map
(42) $\widetilde{\mathrm{qtrs}}:\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P).$
given by $\widetilde{\mathrm{qtrs}}(w_{1}\otimes\vartheta)=(w_{1}\otimes\mathbbm{1})\cdot\mathrm{qtrs}(\vartheta)$ is the inverse map of $\widetilde{\beta}$ [9]. According to [10] we have
Proposition 3.3.
The following properties hold
  1. (1)
    $[\vartheta]_{1}\,[\vartheta]_{2}=\epsilon(\vartheta)\mathbbm{1}$.
  2. (2)
    $(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\Delta_{\Omega^{\bullet}(P)})\circ\mathrm{qtrs}=(\mathrm{qtrs}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\Delta$.
  3. (3)
    $(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{\Omega^{\bullet}(P)})\circ\mathrm{qtrs}=(\sigma\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{\Omega^{\bullet}(P)})\circ(S\otimes\mathrm{qtrs})\circ\Delta$, where
    $\sigma:\Gamma^{\wedge}\otimes\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}$
    is the canonical graded twist map, i.e., $\sigma(\vartheta\otimes w)=(-1)^{kl}\,w\otimes\vartheta$ if $w$in $\Omega^{k}(P)$ and $\vartheta$in $\Gamma^{\wedge l}$.
  4. (4)
    $\mu\,\mathrm{qtrs}(\vartheta)=(-1)^{lk}\mathrm{qtrs}(\vartheta)\,\mu$ for all $\mu$in $\Omega^{k}(B)$, $\vartheta$in $\Gamma^{\wedge l}.$
Let
$\mathfrak{f}_{1},\,\mathfrak{f}_{2}:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P)$
be two graded linear maps. The convolution product of $\mathfrak{f}_{1}$ with $\mathfrak{f}_{2}$ is defined by
$\mathfrak{f}_{1}\ast\mathfrak{f}_{2}=m_{\Omega}\circ(\mathfrak{f}_{1}\otimes\mathfrak{f}_{2})\circ\Delta:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P).$
Henceforth we will just consider graded maps $\mathfrak{f}$ such that
(43) $\mathfrak{f}(\mathbbm{1})=\mathbbm{1}\qquad\mbox{and}\qquad(\mathfrak{f}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\mathrm{Ad}=\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{f},$
where $\mathrm{Ad}:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}\otimes\Gamma^{\wedge}$ is the extension of the right $\mathcal{G}$–coaction $\mathrm{Ad}:H\longrightarrow H\otimes H$ (see equation (19)). We say that $\mathfrak{f}$ is a convolution invertible map if there exists a graded linear map
$\mathfrak{f}^{-1}:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P)$
such that
(44) $\mathfrak{f}\ast\mathfrak{f}^{-1}=\mathfrak{f}^{-1}\ast\mathfrak{f}=\mathbbm{1}\epsilon.$
A direct calculation shows that the set of all convolution invertible maps $\{\mathfrak{f}:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P)\}$ is a group with respect to the convolution product. The next proposition is one of the purposes of this paper
Proposition 3.4.
There exist a group isomorphism between the group of all convolution invertible maps and the group of all graded left $\Omega^{\bullet}(B)$–module isomorphisms
$\mathfrak{F}:\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)$
that satisfy
(45) $\mathfrak{F}(\mathbbm{1})=\mathbbm{1}\qquad\mbox{ and }\qquad(\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\Delta_{\Omega^{\bullet}(P)}=\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{F}.$
Here we are considering the group product $(\mathfrak{F}_{1}\circ\mathfrak{F}_{2})(w)=\mathfrak{F}_{2}(\mathfrak{F}_{1}(w))$.
Proof.
Let us start by considering a map $\mathfrak{F}$. Then we define a graded linear map
(46) $\mathfrak{f}_{\mathfrak{F}}:=m_{\Omega^{\bullet}}\circ(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F})\circ\mathrm{qtrs}:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P),$
where $m_{\Omega^{\bullet}}:\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)$ is the product map. We are going to show that $\mathfrak{f}_{\mathfrak{F}}$ is a convolution invertible map. First, since $\mathrm{qtrs}(\mathbbm{1})=\mathbbm{1}\otimes_{B}\mathbbm{1}$ it follows $\mathfrak{f}_{\mathfrak{F}}(\mathbbm{1})=\mathbbm{1}.$ Secondly, according to Proposition 3.3 point $3$
$\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{f}_{\mathfrak{F}}$ $=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}(\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{F}))\circ\mathrm{qtrs}$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\mathrm{id}_{\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}}\otimes_{\Omega^{\bullet}(B)}(\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{F}))\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{\Omega^{\bullet}(P)})\circ\mathrm{qtrs}$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\mathrm{id}_{\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}}\otimes_{\Omega^{\bullet}(B)}(\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{F}))\circ(\sigma\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{\Omega^{\bullet}(P)})\circ(S\otimes\mathrm{qtrs})\circ\Delta,$
where
$\widehat{m}_{\Omega^{\bullet}}:(\Omega^{\bullet}(P)\otimes\Gamma^{\wedge})\otimes_{\Omega^{\bullet}(B)}(\Omega^{\bullet}(P)\otimes\Gamma^{\wedge})\longrightarrow\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}$
is such that $\widehat{m}_{\Omega^{\bullet}}(w_{1}\otimes\vartheta_{1}\otimes_{\Omega^{\bullet}(B)}w_{2}\otimes\vartheta_{2})=(-1)^{kl}w_{1}w_{2}\otimes\vartheta_{1}\vartheta_{2}$ if $w_{2}$in $\Omega^{k}(P)$ and $\vartheta_{1}$in $\Gamma^{\wedge l}$. On the other hand, by equation (45) and Proposition 3.3 point $2$
$\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{f}_{\mathfrak{F}}$ $=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}(\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{F}))\circ\mathrm{qtrs}$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}((\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\Delta_{\Omega^{\bullet}(P)}))\circ\mathrm{qtrs}$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}(\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}}))\circ(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\Delta_{\Omega^{\bullet}(P)})\circ\mathrm{qtrs}$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}(\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}}))\circ(\mathrm{qtrs}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\Delta;$
but by considering again Proposition 3.3 point $3$
$\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{f}_{\mathfrak{F}}$ $=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}(\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}}))\circ(\mathrm{qtrs}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\Delta$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\mathrm{id}_{\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}}\otimes_{\Omega^{\bullet}(B)}(\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}}))\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}})$
$\circ\,(\mathrm{qtrs}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\Delta$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\mathrm{id}_{\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}}\otimes_{\Omega^{\bullet}(B)}(\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}}))$
$\circ\,[(({\sigma}\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{\Omega^{\bullet}(P)})\circ(S\otimes\mathrm{qtrs})\circ\Delta)\otimes\mathrm{id}_{\Gamma^{\wedge}}]\circ\Delta.$
So for all $\vartheta$in $\Gamma^{\wedge}$
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{f}_{\mathfrak{F}}(\vartheta))$ $=$ $(-1)^{\partial\vartheta^{(1)}(\partial[\vartheta^{(2)}]_{1}+\partial[\vartheta^{(2)}]_{2})}\,[\vartheta^{(2)}]_{1}F([\vartheta^{(2)}]_{2})\otimes S(\vartheta^{(1)})\vartheta^{(3)}$
$=$ $(-1)^{\partial\vartheta^{(1)}\partial\vartheta^{(2)}}\,[\vartheta^{(2)}]_{1}F([\vartheta^{(2)}]_{2})\otimes S(\vartheta^{(1)})\vartheta^{(3)}\;=\;(\mathfrak{f}_{\mathfrak{F}}\otimes\mathrm{id}_{\Gamma^{\wedge}})\mathrm{Ad}(\vartheta),$
because $\mathrm{qtrs}$ is a graded linear map. Finally, consider
$\mathfrak{f}_{\mathfrak{F}^{-1}}:=m_{\Omega^{\bullet}}\circ(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\circ\mathrm{qtrs}:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P).$
Then for all $\vartheta$in $\Gamma^{\wedge}$
$(\mathfrak{f}_{\mathfrak{F}}\ast\mathfrak{f}_{\mathfrak{F}^{-1}})(\vartheta)=[\vartheta^{(1)}]_{1}\,\underbrace{\mathfrak{F}([\vartheta^{(1)}]_{2})\,[\vartheta^{(2)}]_{1}}\,\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2}).$
We claim that the expression in the brace is an element of $\Omega^{\bullet}(B)$. Indeed, notice that
(49) $\Delta_{\Omega^{\bullet}(P)}((\mathfrak{f}_{\mathfrak{F}}\ast\mathfrak{f}_{\mathfrak{F}^{-1}})(\vartheta))$ $=$ $\Delta_{\Omega^{\bullet}(P)}\left([\vartheta^{(1)}]_{1}\,\mathfrak{F}([\vartheta^{(1)}]_{2})\,[\vartheta^{(2)}]_{1}\,\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2})\right)$
$=$ $\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([\vartheta^{(1)}]_{2})\,[\vartheta^{(2)}]_{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2})).$
Also we know
$\Delta_{\Omega^{\bullet}(P)}((\mathfrak{f}_{\mathfrak{F}}\ast\mathfrak{f}_{\mathfrak{F}^{-1}})(\vartheta))=\Delta_{\Omega^{\bullet}(P)}(\mathfrak{f}_{\mathfrak{F}}(\vartheta^{(1)}))\cdot\Delta_{\Omega^{\bullet}(P)}(\mathfrak{f}_{\mathfrak{F}^{-1}}(\vartheta^{(2)})).$
By applying equation (3) on $\Delta_{\Omega^{\bullet}(P)}(\mathfrak{f}_{\mathfrak{F}}(\vartheta^{(1)}))$ and equation (3) on $\Delta_{\Omega^{\bullet}(P)}(\mathfrak{f}_{\mathfrak{F}}(\vartheta^{(2)}))$ we have
$\Delta_{\Omega^{\bullet}(P)}((\mathfrak{f}_{\mathfrak{F}}\ast\mathfrak{f}_{\mathfrak{F}^{-1}})(\vartheta))$ $=$ $(-1)^{\partial\vartheta^{(3)}\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\vartheta^{(2)})$
$([\vartheta^{(4)}]_{1}\otimes S(\vartheta^{(3)}))\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(4)}]_{2}))$
$=$ $(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(4)}]_{1}\otimes\vartheta^{(2)}S(\vartheta^{(3)}))$
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(4)}]_{2})).$
On the other hand, by the definition of $\epsilon$ (see equation (18)) and the graded Hopf $\ast$–algebra structure of $\Gamma^{\wedge\,\infty}$ and we get
$\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(2)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2}))$
$=\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\mathbbm{1})\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(2)}))$
[…]
$=(-1)^{\partial\vartheta^{(2)}\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\;\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\underbrace{\epsilon(\vartheta^{(2)})\mathbbm{1}}_{0-\mathrm{degree}})\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(3)})).$
$=(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\underbrace{\vartheta^{(2)}S(\vartheta^{(3)}}_{0-\mathrm{degree}}))\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(4)}))$
$=(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\underbrace{\vartheta^{(2)}S(\vartheta^{(3)}}_{0-\mathrm{degree}}))\,([\vartheta^{(4)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(4)}]_{2}))$
$=(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(4)}]_{1}\otimes\vartheta^{(2)}S(\vartheta^{(3)}))\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(4)}]_{2}))$
and hence
$\Delta_{\Omega^{\bullet}(P)}((\mathfrak{f}_{\mathfrak{F}}\ast\mathfrak{f}_{\mathfrak{F}^{-1}})(\vartheta))=\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(2)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2})).$
By equation (49) and the last equality we conclude
$\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([\vartheta^{(1)}]_{2})\,[\vartheta^{(2)}]_{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2}))$
$=\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(2)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2})),$
which proves our claim. Since $\mathfrak{F}^{-1}$ is a left $\Omega^{\bullet}(B)$–module morphism, by our previous claim and Proposition 3.3 point $1$
$(\mathfrak{f}_{\mathfrak{F}}\ast\mathfrak{f}_{\mathfrak{F}^{-1}})(\vartheta)\;=\;[\vartheta^{(1)}]_{1}\,\mathfrak{F}([\vartheta^{(1)}]_{2})\,[\vartheta^{(2)}]_{1}\,\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2})$ $=$ $[\vartheta^{(1)}]_{1}\,\mathfrak{F}^{-1}(\mathfrak{F}([\vartheta^{(1)}]_{2})\,[\vartheta^{(2)}]_{1}[\vartheta^{(2)}]_{2})$
$=$ $[\vartheta^{(1)}]_{1}\,\mathfrak{F}^{-1}(\mathfrak{F}([\vartheta^{(1)}]_{2})\,\epsilon(\vartheta^{(2)}))$
$=$ $\epsilon(\vartheta^{(1)})\epsilon(\vartheta^{(2)})\mathbbm{1}$
$=$ $\epsilon(\vartheta)\mathbbm{1}$
for every $\vartheta$in $\Gamma^{\wedge}$. In a similar way, we can prove that $\mathfrak{f}_{\mathfrak{F}^{-1}}\ast\mathfrak{f}_{\mathfrak{F}}=\mathbbm{1}\epsilon$ and hence $\mathfrak{f}_{\mathfrak{F}}$ is a convolution invertible map.
Conversely, for a given convolution invertible map $\mathfrak{f}$ let us define the graded linear map
(50) $\mathfrak{F}_{\mathfrak{f}}:=m_{\Omega}\circ(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})\circ\Delta_{\Omega^{\bullet}(P)}:\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P).$
We are going to prove that $\mathfrak{F}_{\mathfrak{f}}$ is a graded left $\Omega^{\bullet}(B)$–module isomorphism which satisfies equation (45). First of all, by equation (43) it is obvious that $\mathfrak{F}_{\mathfrak{f}}(\mathbbm{1})=\mathbbm{1}$. Secondly, taking $\mu$in $\Omega^{\bullet}(B)$ and $w$in $\Omega^{\bullet}(P)$ we have $\mathfrak{F}_{\mathfrak{f}}(\mu w)=m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})(\mu\otimes\mathbbm{1})\Delta_{\Omega^{\bullet}(P)}(w)=\mu\,\mathfrak{F}_{f}(w).$ Thirdly, by equation (43)
$\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{F}_{\mathfrak{f}}$ $=$ $\widehat{m}_{\Omega}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes(\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{f}))\circ\Delta_{\Omega^{\bullet}(P)}$
$=$ $\widehat{m}_{\Omega}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes((\mathfrak{f}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\mathrm{Ad}))\circ\Delta_{\Omega^{\bullet}(P)},$
where
$\widehat{m}_{\Omega}:(\Omega^{\bullet}(P)\otimes\Gamma^{\wedge})\otimes(\Omega^{\bullet}(P)\otimes\Gamma^{\wedge})\longrightarrow\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}$
[…]
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}_{\mathfrak{f}}(w))$ $=$ $(-1)^{\partial w^{3}(\partial w^{1}+\partial w^{2})}w^{(0)}\mathfrak{f}(w^{(3)})\otimes w^{(1)}S(w^{(2)})w^{(4)}$
$=$ $(-1)^{\partial w^{3}\partial w^{1}}w^{(0)}\mathfrak{f}(w^{(2)})\otimes\epsilon(w^{(1)})w^{(3)}$
$=$ $w^{(0)}\mathfrak{f}(w^{(2)})\otimes\epsilon(w^{(1)})w^{(3)}$
$=$ $w^{(0)}\mathfrak{f}(w^{(2)})\otimes\epsilon(w^{(1)})w^{(3)}$
$=$ $w^{(0)}\mathfrak{f}(w^{(1)})\otimes w^{(2)}=(\mathfrak{F}_{\mathfrak{f}}\otimes\mathrm{id}_{\Gamma^{\wedge}})\Delta_{\Omega^{\bullet}(P)}(w).$
Finally, notice that for all $w$in $\Omega^{\bullet}(P)$
$\mathfrak{F}_{\mathfrak{f}^{-1}}(\mathfrak{F}_{\mathfrak{f}}(w))\;=\;m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f}^{-1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}_{\mathfrak{f}}(w))$ $=$ $m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f}^{-1})(\mathfrak{F}_{\mathfrak{f}}\otimes\mathrm{id}_{\Gamma^{\wedge}})\,\Delta_{\Omega^{\bullet}(P)}(w)$
$=$ $\mathfrak{F}_{f}(w^{(0)})\,\mathfrak{f}^{-1}(w^{(1)})$
$=$ $w^{(0)}\,\mathfrak{f}(w^{(1)})\,\mathfrak{f}^{-1}(w^{(2)})$
$=$ $w^{(0)}\,\epsilon(w^{(1)})\;=\;w.$
A similar calculation shows $\mathfrak{F}_{\mathfrak{f}}(\mathfrak{F}_{\mathfrak{f}^{-1}}(w))=w$, so $\mathfrak{F}_{\mathfrak{f}}$ is biyective and $\mathfrak{F}^{-1}_{\mathfrak{f}}=\mathfrak{F}_{\mathfrak{f}^{-1}}.$
Our next step is to prove that
$\mathfrak{F}\xmapsto{\widehat{\Lambda}}\mathfrak{f}_{\mathfrak{F}}\,,\qquad\mathfrak{f}\xmapsto{\widetilde{\Lambda}}\mathfrak{F}_{\mathfrak{f}}$
are mutually inverse. Notice that for all $w$in $\Omega^{\bullet}(P)$
$\mathfrak{F}_{\mathfrak{f}_{\mathfrak{F}}}(w)=m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes(m_{\Omega^{\bullet}}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F})\mathrm{qtrs}))\Delta_{\Omega^{\bullet}(P)}(w)=\underbrace{w^{(0)}[w^{(1)}]_{1}}\mathfrak{F}([w^{(1)}]_{2}),$
where the expression in the brace is an element of $\Omega^{\bullet}(B)$. In fact,
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}_{\mathfrak{f}_{\mathfrak{F}}}(w))=\Delta_{\Omega^{\bullet}(P)}\left(w^{(0)}[w^{(1)}]_{1}\mathfrak{F}([w^{(1)}]_{2})\right)=\Delta_{\Omega^{\bullet}(P)}(w^{(0)}[w^{(1)}]_{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(1)}]_{2}));$
however by equation (3) we get
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}_{\mathfrak{f}_{\mathfrak{F}}}(w))\;=\;\Delta_{\Omega^{\bullet}(P)}(w^{(0)}\mathfrak{f}_{\mathfrak{F}}(w^{(1)}))$ $=$ $(w^{(0)}\otimes w^{(1)})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{f}_{\mathfrak{F}}(w^{(2)}))$
$=$ $(-1)^{\partial w^{(2)}\partial[w^{(3)}]_{1}}(w^{(0)}\otimes w^{(1)})$
$([w^{(3)}]_{1}\otimes S(w^{(2)}))\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(3)}]_{2}))$
$=$ $(-1)^{(\partial w^{(1)}+\partial w^{(2)})\partial[w^{(3)}]_{1}}\,(w^{(0)}[w^{(3)}]_{1}\otimes w^{(1)}S(w^{(2)}))$
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(3)}]_{2}))$
[…]
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(2)}]_{2}))$
$=$ $(w^{(0)}[w^{(2)}]_{1}\otimes\epsilon(w^{(1)})\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(2)}]_{2}))$
$=$ $(w^{(0)}\otimes\epsilon(w^{(1)})\mathbbm{1})([w^{(2)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(2)}]_{2}))$
$=$ $(w^{(0)}\otimes\epsilon(w^{(1)})\mathbbm{1})\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F})\mathrm{qtrs}(w^{(2)}))$
$=$ $(w^{(0)}\otimes\mathbbm{1})\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F})\mathrm{qtrs}(w^{(1)}))$
$=$ $(w^{(0)}[w^{(1)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(1)}]_{2})).$
This implies that
$\Delta_{\Omega^{\bullet}(P)}(w^{(0)}[w^{(1)}]_{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(1)}]_{2}))=(w^{(0)}[w^{(1)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(1)}]_{2})),$
which proves our assertion. Since $\mathfrak{F}$ is a left $\Omega^{\bullet}(B)$–module morphism, by this assertion and Proposition 3.3 point $1$, we have
$\mathfrak{F}_{\mathfrak{f}_{\mathfrak{F}}}(w)=w^{(0)}[w^{(1)}]_{1}\,\mathfrak{F}([w^{(1)}]_{2})=\mathfrak{F}(w^{(0)}[w^{(1)}]_{1}[w^{(1)}]_{2}))=\mathfrak{F}(w^{(0)}\epsilon(w^{(1)}))=\mathfrak{F}(w).$
On the other hand, for every $\vartheta$in $\Gamma^{\wedge}$
$\mathfrak{f}_{\mathfrak{F}_{\mathfrak{f}}}(\vartheta)$ $=$ $m_{\Omega^{\bullet}}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}(m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})\Delta_{\Omega^{\bullet}(P)}))\mathrm{qtrs}(\vartheta)$
$=$ $[\vartheta]_{1}[\vartheta]^{(0)}_{2}\mathfrak{f}([\vartheta]^{(1)}_{2})$
$=$ $m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})([\vartheta]_{1}\otimes\mathbbm{1})([\vartheta]^{(0)}_{2}\otimes[\vartheta]^{(1)}_{2})$
$=$ $m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})([\vartheta]_{1}\otimes\mathbbm{1})\Delta_{\Omega^{\bullet}(P)}([\vartheta]_{2})$
$=$ $m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})\,\widetilde{\beta}(\mathrm{qtrs}(\vartheta))$
$=$ $m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})(\mathbbm{1}\otimes\vartheta)\;=\;\mathfrak{f}(\vartheta)$
and hence $\widetilde{\Lambda}=\widehat{\Lambda}^{-1}$. Finally, to complete the proof it is enough to prove that $\widehat{\Lambda}$ or $\widehat{\Lambda}^{-1}$ is a group morphism. In fact
$\mathfrak{F}_{\mathfrak{f}_{1}\ast\mathfrak{f}_{2}}(w)=m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes(\mathfrak{f}_{1}\ast\mathfrak{f}_{2}))\Delta_{\Omega^{\bullet}(P)}(w)=w^{(0)}\,\mathfrak{f}_{1}(w^{(1)})\,\mathfrak{f}_{2}(w^{(2)});$
while by equation (45)
$\mathfrak{F}_{\mathfrak{f}_{2}}(\mathfrak{F}_{\mathfrak{f}_{1}}(w))\;=\;m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f}_{2})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}_{\mathfrak{f}_{1}}(w))$ $=$ $m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f}_{2})(\mathfrak{F}_{\mathfrak{f}_{1}}\otimes\mathrm{id}_{\Gamma^{\wedge}})\,\Delta_{\Omega^{\bullet}(P)}(w)$
$=$ $\mathfrak{F}_{\mathfrak{f}_{1}}(w^{(0)})\,\mathfrak{f}_{2}(w^{(1)})$
$=$ $w^{(0)}\,\mathfrak{f}_{1}(w^{(1)})\,\mathfrak{f}_{2}(w^{(2)})$
for all $w$in $\Omega^{\bullet}(P)$. Therefore $\mathfrak{f}_{1}\ast\mathfrak{f}_{2}\xmapsto{\widehat{\Lambda}^{-1}}\mathfrak{F}_{\mathfrak{f}_{1}}\circ\mathfrak{F}_{\mathfrak{f}_{2}}$ and the proposition follows. ∎
4. The form of the maps $\{T^{\mathrm{L}}_{k}\}$
Let $G$ be a compact matrix Lie group, and let $\mathcal{G}$ be its associated quantum group [2]. If $\delta^{V}$ is a (unitary) irreducible $\mathcal{G}$–corepresentation, then it induces a $G$–representation on $V$. Indeed, by Theorem 1.1 we know that
$\delta^{V}(e_{j})=\sum^{n}_{i=1}e_{i}\otimes g^{V}_{ij}.$
Then the linear map
$\alpha^{V}:G\times V\longrightarrow V$
given by
$\alpha^{V}(A,e_{j})=\sum^{n}_{i=1}g^{V}_{ij}(A)\,e_{i}$
is a unitary and irreducible $G$–representation. The following proposition is another of our purposes.
Proposition 4.1.
Let $G$ be a compact matrix Lie group and let $\pi:P\longrightarrow B$ be a classical principal $G$–bundle, where $P$ is the total space, $B$ is the base space and $\pi$ is the bundle projection. Assume $P$, $B$ are compact. If $\mathcal{T}$ is a complete set of mutually non–equivalent irreducible $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$, then for every $\delta^{V}$in $\mathcal{T}$, there exists $\{T^{\mathrm{L}}_{k}\}^{d_{V}}_{k=1}$subseteq $\textsc{Mor}(\delta^{V},\gamma^{\#})$, for some $d_{V}$in $\mathbb{N}$, such that equation (30) holds. Here
$\gamma^{\#}:C^{\infty}_{\mathbb{C}}(P)\longrightarrow C^{\infty}_{\mathbb{C}}(P\times G)\cong C^{\infty}_{\mathbb{C}}(P)\otimes C^{\infty}_{\mathbb{C}}(G)$
is the pull–back of the right $G$–action on $P$. In this setting, $C^{\infty}_{\mathbb{C}}(P)$ denotes the space of $\mathbb{C}$–valued smooth functions of $P$ and the tensor product is taken to be the completed injective tensor product.
Proof.
Let $\delta^{V}$in $\mathcal{T}$ and consider its associated representation $\alpha^{V}$. Since $\alpha^{V}$ is unitary, the associated vector bundle $\pi_{\alpha^{V}}:E^{V}\longrightarrow B$ is a Hermitian bundle.
For each $b$in $B$, let $(U_{b},\Phi_{b})$ be a local trivialization of the Hermitian bundle $\pi_{\alpha^{V}}:E^{V}\longrightarrow B$ around $b$ associated to a principal $G$–bundle local trivialization $(U_{b},\Psi_{b})$ of $\pi:P\longrightarrow B$. Then there exists a set of local sections $\{\hat{s}^{b}_{1},\ldots,\hat{s}^{b}_{1}\}$subseteq $\Gamma\bigl{(}\pi_{{\alpha^{V}}}^{-1}(U_{b})\bigr{)}$ such that $\hat{s}^{b}_{i}(a)=(a,e_{i})$ for all $a$in $U_{b}$ (under the diffeomorphism $\Phi_{b}$). Since $\{U_{b}\}_{b\in B}$ is an open cover, by compactness, there exist points $b_{1},\ldots,b_{r}\in B$ such that $\{\,U_{b_{i}}\}_{i=1}^{r}$ remains an open cover of $B$. Let $\{\rho_{b_{i}}\}_{i=1}^{r}$ be a partition of unity with $\rho_{b_{i}}$ having compact support and subordinate to the open cover $\{\,U_{b_{i}}\}_{i=1}^{r}$. In this way, consider the global sections of $E^{V}$
$\{s^{i}_{j}=\sqrt{\rho_{b_{i}}}\,\hat{s}^{b_{i}}_{j}\}^{r,n}_{i,j=1}.$
For every $x$in $P$, if $a=\pi(x)$in $B$, it is well–known that there exists a unique $v$in $V$ such that $s^{i}_{j}(a)=[x,v]$in $E^{V}$. This induces a smooth map
$f^{i}_{j}:P$ $\longrightarrow V$
$x$ $\longrightarrow v=\sqrt{\rho_{b_{i}}(\pi(x))}\;\alpha^{V}(A(x)^{-1},e_{j}),$
where $A(x)$ is the unique element of $G$ such that $\Psi_{b_{i}}(\pi(x),A(x))=x$. These maps satisfy $f^{i}_{j}(\gamma(x,B))=\alpha^{V}(B^{-1},f^{i}_{j}(x))$ for every $x$in $P$, $B$in $G$. We define the smooth functions
$f^{i}_{jk}:P$ $\longrightarrow\mathbb{C}$
$x$ $\longrightarrow\langle e_{k}|f^{i}_{j}(x)\rangle,$
where $\langle$-|$\rangle$ - | - ⟩ is the inner product that makes $\delta^{V}$ unitary (antilinear in the second coordinate). Now, let us consider the linear maps
$T^{i}_{j}:V\longrightarrow C^{\infty}_{\mathbb{C}}(P)$
given by $T^{i}_{j}(e_{k})=f^{i}_{jk}$. A direct calculation shows that for all $x$in $P$ and for all $A$in $G$
$\left(\sum^{r,n}_{l,k=1}f^{l\,\ast}_{ki}f^{l}_{kj}\right)(x)=\delta_{ij},\qquad(\gamma^{\#}\circ T^{i}_{j})(x,B)=((T^{i}_{j}\otimes\mathrm{id}_{H})\circ\delta^{V})(x,B).$
Then proposition follows by taking $T^{\mathrm{L}}_{1}=T^{1}_{1}$, $T^{\mathrm{L}}_{2}=T^{1}_{2}$,…, $T^{\mathrm{L}}_{n}=T^{1}_{n}$, $T^{\mathrm{L}}_{n+1}=T^{2}_{1}$,…, $T^{\mathrm{L}}_{d_{V}}=T^{r}_{n}$ with $d_{V}=rn$. ∎
4.1. Quantum Principal Bundles and Dunkl Operators
For an example of the theory presented in [1], we will use somewhat lesser-known quantum bundles: the quantum principal $\mathcal{G}$–bundles developed in [6, 11]. We shall denote them by
$\zeta=(P,B,\Delta_{P}).$
These qpb’s are defined by the dualization of classical principal bundles with (finite) Coxeter groups $W(H)$ as the structure group, which we shall denote them by
$\rho:P_{\mathrm{class}}\longrightarrow P_{\mathrm{class}}/W(H),$
where $P_{\mathrm{class}}$ is the classical total space, $P_{\mathrm{class}}/W(H)$ is the classical base space, and the map $\rho:P_{\mathrm{class}}\longrightarrow P_{\mathrm{class}}/W(H)$ is the canonical projection (which is the bundle projection).
On the other hand, horizontal forms on $\zeta$ are given by the complexification of the de–Rham graded differential $\ast$–algebra of $P_{\mathrm{class}}$, quantum differential forms of $B$ are given by the complexification of the de–Rham graded differential $\ast$–algebra of $P_{\mathrm{class}}/W(H)$, and $(\Gamma,d)$ is given by the theory of $\ast$–FODC’s on finite groups [11]. It is worth mentioning that in $\zeta$ there exists a canonical qpc
(51) $\omega^{c}:\mathfrak{qg}^{\#}$ $\longrightarrow\Omega^{1}(P)$
$\theta$ $\longmapsto\mathbbm{1}_{B}\otimes\theta.$
and its covariant derivative $D^{\omega^{c}}$ is exactly the de-Rham differential. Since $\omega^{c}$ is real and regular, $\widehat{D}^{\omega^{c}}=D^{\omega^{c}}$.
In order to give a concrete example let us focus our study on standard Dunkl connections using the development in [11]. It is worth mentioning that, since we have changed the standard definition of qpc’s in order to embrace a more general theory, for us Dunkl displacements $\lambda:\mathfrak{qg}^{\#}\longrightarrow\Omega^{1}(P)$ do not need the $i=\sqrt{-1}$ factor. These qpc’s are given by
$\omega=\omega^{\mathrm{c}}+\lambda,$
where $\lambda$ satisfies $\lambda(\pi(\phi_{\sigma_{r}}))=\varrho_{r}r^{\#}$,
$\varrho_{r}:P_{\mathrm{class}}$ $\longrightarrow\mathbb{R}$
$x$ $\longrightarrow{\kappa(r)\over\langle r|x\rangle},$
$\kappa:R\longrightarrow\mathbb{R}$ is a multiplicative function (it is $W(H)$–invariant), $r$in $R$ with $R$ the corresponding root system, and $r^{\#}$ is the element of the dual space associated with $r$. Covariant derivatives of this kind of qpc’s are given by
$(D^{\omega}f)(x)=df(x)+\sum_{r\in R^{+}}\kappa(r){f(x)-f(x\sigma_{r})\over\langle r|x\rangle}r$
for every $f$in $P$, which is a Dunkl operator in vector form [6].
Let $\delta^{V}$in $\mathcal{T}$. Then
$\nabla^{\omega}_{V}T=\sum^{n_{V}}_{i=1}\mu^{D^{\omega}T}_{i}\otimes_{B}T^{V}_{i}\,,\quad\widehat{\nabla}^{\omega}_{V}T=\sum^{n_{V}}_{i=1}T^{V}_{i}\otimes_{B}\mu^{D^{\omega}T}_{i},$
where $\mu^{D^{\omega}T}_{k}=\displaystyle\sum^{n_{V}}_{i=1}D^{\omega}T(e_{i})f^{V\,\ast}_{ki}$ and $T(e_{k})$in $P$ [1]. Notice that Dunkl connections are multiplicative, but not regular [6]. Furthermore, it is worth mentioning that, in differential geometry, there are no principal connections on the principal bundle $\rho:P_{\mathrm{class}}\longrightarrow P_{\mathrm{class}}/W(H)$.
On the other hand, the left and right canonical Hermitian structures are given by
$\langle T_{1},T_{2}\rangle_{\mathrm{L}}=\sum^{n_{V}}_{k=1}T_{1}(e_{k})T_{2}(e_{k})^{\ast}\,,\qquad\langle T_{1},T_{2}\rangle_{\mathrm{R}}=\sum^{n_{V}}_{k=1}T_{1}(e_{k})^{\ast}T_{2}(e_{k}).$
For an explicit example of Theorem $3.14$ of [1], let us take a real Dunkl connection. These connections are characterized by $\omega=\omega^{\mathrm{c}}+\widetilde{\lambda}$, with $\widetilde{\lambda}=i\lambda$. Therefore, using equation (30),
$\langle\nabla^{\omega}_{V}(T_{1}),T_{2}\rangle_{\mathrm{L}}\,=\,\sum^{n_{V}}_{i=1}\mu^{D^{\omega}T_{1}}_{i}\langle T^{V}_{i},T_{2}\rangle_{\mathrm{L}}\,=\,\sum^{n_{V}}_{k=1}\left(D^{\omega}T_{1}(e_{k})\right)T_{2}(e_{k})^{\ast};$
and evaluating on any $x$in $P_{\mathrm{class}}$ we have
$\sum^{n_{V}}_{k=1}\left(dT_{1}(e_{k})(x)T_{2}(e_{k})^{\ast}(x)+i\sum_{r\in R^{+}}\kappa(r){T_{1}(e_{k})(x)-T_{1}(e_{k})(x\sigma_{r})\over\langle r|x\rangle}r\,T_{2}(e_{k})^{\ast}(x)\right).$
In the same way
$\langle T_{1},\nabla^{\omega}_{V}(T_{2})\rangle_{\mathrm{L}}\,=\,\sum^{n_{V}}_{i=1}\langle T_{1},T_{i}\rangle_{\mathrm{L}}\left(\mu^{D^{\omega}T_{2}}_{i}\right)^{\ast}\,=\,\sum^{n_{V}}_{k=1}T_{1}(e_{k})\left(D^{\omega}T_{2}(e_{k})\right)^{\ast};$
and evaluating on any $x$in $P_{\mathrm{class}}$ we get
$\sum^{n_{V}}_{k=1}\left(T_{1}(e_{k})(x)dT_{2}(e_{k})^{\ast}(x)-i\,T_{1}(e_{k})(x)\sum_{r\in R^{+}}\kappa(r){T_{2}(e_{k})^{\ast}(x)-T_{2}(e_{k})^{\ast}(x\sigma_{r})\over\langle r|x\rangle}r\right).$
This implies for all $x$in $P_{\mathrm{class}}$ that $\langle\nabla^{\omega}_{V}(T_{1}),T_{2}\rangle_{\mathrm{L}}+\langle T_{1},\nabla^{\omega}_{V}(T_{2})\rangle_{\mathrm{L}}$ is equal to
$d\langle T_{1},T_{2}\rangle_{\mathrm{L}}(x)+i\sum_{r\in R^{+}}\kappa(r){q(x)-q(x\sigma_{r})\over\langle r|x\rangle}r,$
where $q=\displaystyle\sum^{n_{V}}_{k=1}T_{1}(e_{k})(\sigma_{r}\cdot T_{2}(e_{k}))^{\ast}$in $P$. However, since $T_{1}$, $T_{2}$in $\textsc{Mor}(\delta^{V},\Delta_{P})$, we have $\sigma_{r}\cdot T_{i}(e_{j})=\displaystyle\sum^{n_{V}}_{k=1}\lambda^{\sigma_{r}}_{kj}T_{i}(e_{k})$. Thus $q=\displaystyle\sum^{n_{V}}_{k=1}(\sigma_{r}\cdot T_{1}(e_{k}))T_{2}(e_{k})^{\ast}$ and hence $q(x)=q(x\sigma_{r})$ for all $x$in $P_{\mathrm{class}}$. In summary, we have proven explicitly that
$\langle\nabla^{\omega}_{V}(T_{1}),T_{2}\rangle_{\mathrm{L}}+\langle T_{1},\nabla^{\omega}_{V}(T_{2})\rangle_{\mathrm{L}}=d\langle T_{1},T_{2}\rangle_{\mathrm{L}}.$
A similar argument proves that
^$\langle\widehat{\nabla}^{\omega}_{V}(T_{1}),T_{2}\rangle_{\mathrm{R}}+\langle T_{1},\widehat{\nabla}^{\omega}_{V}(T_{2})\rangle_{\mathrm{R}}=d\langle T_{1},T_{2}\rangle_{\mathrm{R}}.$
Defining $g^{V}_{ij}:W(H)\longrightarrow\mathbb{C}$ by $g^{V}_{ij}(h)=\lambda^{h}_{ij}$ we get
$\mathrm{qtrs}(g^{V}_{ij})=\sum^{n_{V}}_{k=1}f^{V\,\ast}_{ki}\otimes_{B}f^{V}_{kj},$
and by considering $\omega^{c}$ in equation (37), we have for all $\theta$in $\mathfrak{qg}^{\#}$
(52) $\mathrm{qtrs}(\theta)=\mathbbm{1}\otimes_{\Omega^{\bullet}(B)}(\mathbbm{1}\otimes\theta)-(\mathbbm{1}\otimes\theta^{(0)})\,\mathrm{qtrs}(\theta^{(1)}),$
where $\mathrm{ad}(\theta)=\theta^{(0)}\otimes\theta^{(1)}$.
Now let us consider a graded left $\Omega^{\bullet}(B)$–module isomorphism
$\mathfrak{F}:\Omega^{\bullet}(B)\longrightarrow\Omega^{\bullet}(B)$
which (for irreducible degree $1$ elements) is given by
$\mathfrak{F}(\mu+x\otimes\theta)=\mu+x\lambda(\theta)+x\otimes\theta,$
where $\lambda$ is a Dunkl displacement, $\mu$in $\mathrm{Hor}^{1}P$, $x$in $P$, and $\theta$in $\mathfrak{qg}^{\#}$. This map is actually a qgt $\mathfrak{f}$ and a direct calculation shows that $\mathfrak{F}^{\circledast}\omega^{c}$ is a standard Dunkl connection. By using the form of the last map $\mathfrak{F}$, every single classical gauge transformation can be extended into a quantum one.