### v2

# Some Constructions on Quantum Principal Bundles

Abstract.
This paper works as an appendix of the paper titled Geometry of Associated Quantum Vector Bundles and the Quantum Gauge Group. Here, we are going to prove four statements in the theory of quantum principal bundles:
  1. (1)
    The universal differential envelope $\ast$–calculus of a matrix (compact) Lie group, for the classical bicovariant First Order Differential $\ast$–Calculus, is the algebra of differential forms.
  2. (2)
    An example of a quantum principal bundle in which the space of base forms is not generated by the base space.
  3. (3)
    The group isomorphism between convolution-invertible maps and covariant left module isomorphisms at the level of differential calculus
  4. (4)
    The way the maps $\{T^{V}_{k}\}$ from Remark 3.1 look in differential geometry.
Furthermore, we are going to present an example of the theory of the reference [1] in a special classical/quantum hybrid principal bundle.
1. Universal Differential Envelope $\ast$–calculus in Differential Geometry
A compact matrix quantum group (or a just a quantum group) will be denoted by $\mathcal{G}$, and its dense $\ast$–Hopf (sub)algebra will be denoted by
(1) $H^{\infty}:=(H,\cdot,\mathbbm{1},\Delta,\epsilon,S,\ast),$
where $\Delta$ is the coproduct, $\epsilon$ is the counity and $S$ is the coinverse. The space $H^{\infty}$ shall be treated as the algebra of all polynomial functions defined on $\mathcal{G}$. In the same way a (smooth right) $\mathcal{G}$–corepresentation on a $\mathbb{C}$–vector space $V$ is a linear map
$\delta^{V}:V\longrightarrow V\otimes H$
such that
(2) $(\mathrm{id}_{V}\otimes\epsilon)\circ\delta^{V}\cong\mathrm{id}_{V}$
and
(3) $(\mathrm{id}_{V}\otimes\Delta)\circ\delta^{V}=(\delta^{V}\otimes\mathrm{id}_{H})\circ\delta^{V}.$
We say that the corepresentation is finite–dimensional if $\dim_{\mathbb{C}}(V)<|\mathbb{N}|$. $\delta^{V}$ usually receives the name of the (right) coaction of $\mathcal{G}$ on $V$. It is worth mentioning that in the general theory ([2]), equation (2) is not necessary.
The set of all corepresentation morphisms between two corepresentations $\delta^{V}$, $\delta^{W}$ will be denoted as
(4) $\textsc{Mor}(\delta^{V},\delta^{W}),$
A $\mathcal{G}$–corepresentation $\delta^{V}$ is reducible if there exists a non–trivial subspace $L$ $(L\not=\{0\},V)$ such that $\delta^{V}(L)=L\otimes H$ and $\delta^{V}$ is unitary if viewed as an element of $B(V)\otimes H$ (with $B(V):=\{f:V\longrightarrow V\mid f\mbox{ is linear}\}$) is unitary. Of course, for the last definition, it is necessary an inner product $\langle$-|$\rangle$ - | - ⟩ on $V$. In [2], Woronowicz proved that for every finite–dimensional $\mathcal{G}$–corepresentation on $V$ there is an inner product $\langle$-|$\rangle$ - | - ⟩ (not necessarily unique) making the corepresentation unitary. Thus, henceforth we are going to consider that every finite–dimensional $\mathcal{G}$–corepresentation is unitary. In [2] one can find a proof of the following theorem.
Theorem 1.1.
Let $\mathcal{T}$ be a complete set of mutually non–equivalent irreducible unitary (necessarily finite–dimensional) $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$ (the trivial corepresentation on $\mathbb{C}$). For any $\delta^{V}$in $\mathcal{T}$ that coacts on $(V,\langle-|-\rangle)$,
(5) $\delta^{V}(e_{i})=\sum^{n_{V}}_{j=1}e_{j}\otimes g^{V}_{ji},$
where $\{e_{i}\}^{n_{V}}_{i=1}$ is an ortonormal basis of $V$ and $\{g^{V}_{ij}\}^{n_{V}}_{i,j=1}$subseteq $H$. Then $\{g^{V}_{ij}\}_{\delta^{V},i,j}$ is a linear basis of $H$, where the index $\delta^{V}$ runs on $\mathcal{T}$ and $i$, $j$ run from $1$ to $n_{V}=\mathrm{dim}_{\mathbb{C}}(V)$.
For every $\delta^{V}$in $\mathcal{T}$, the set $\{g^{V}_{ij}\}^{n_{V}}_{ij=1}$ satisfies
(6) $\phi(g^{\alpha}_{ij})=\sum^{n_{\alpha}}_{k=1}g^{\alpha}_{ik}\otimes g^{\alpha}_{kj},\quad\kappa(g^{\alpha}_{ij})=g^{\alpha\,\ast}_{ji},\quad\sum^{n_{\alpha}}_{k=1}g^{\alpha}_{ik}g^{\alpha\,\ast}_{jk}=\sum^{n_{\alpha}}_{k=1}g^{\alpha\,\ast}_{ki}g^{\alpha}_{kj}=\delta_{ij}\mathbbm{1},\quad\epsilon(g^{\alpha}_{ij})=\delta_{ij}$
with $\delta_{ij}$ being the Kronecker delta, among other properties [2].
Let $(\Gamma,d)$ be a bicovariant First Order Differential $\ast$–Calculus ($\ast$–FODC) over $\mathcal{G}$ ([3, 6]) and consider the $\mathbb{C}$–vector space given by
(7) $\mathfrak{qg}^{\#}:=\{\theta\in\Gamma\mid\Phi_{\Gamma}(\theta)=\mathbbm{1}\otimes\theta\}=\mathbbm{1}\otimes{\mathrm{Ker}(\epsilon)\over\mathcal{R}}\cong{\mathrm{Ker}(\epsilon)\over\mathcal{R}}.$
This space allows to consider the quantum germs map
(8) $\pi:H$ $\longrightarrow\mathfrak{qg}^{\#}$
$g$ $\longmapsto[g-\epsilon(g)\mathbbm{1}]_{\mathcal{R}},$
where $[h]_{\mathcal{R}}$ denotes the equivalence class of the element $h$in $H$. The map $\pi$ has several useful properties, for example the restriction map $\pi|_{\mathrm{Ker}(\epsilon)}$ is surjective and
(9) $\ker(\pi)=\mathcal{R}\oplus\mathbb{C}\mathbbm{1},$ $\qquad\;\;\,\pi(g)=S(g^{1})dg^{2},$ $\quad dg=g^{(1)}\pi(g^{(2)}),$
$\pi(g)=-(dS(g^{(1)}))g^{(2)},$ $\quad dS(g)=-\pi(g^{(1)})S(g^{(2)}),$ $\quad\pi(g)^{\ast}=-\pi(S(g)^{\ast})$
for all $g$in $H$ [6]. It is worth mentioning that the canonical right action ${{}_{\Gamma}}\Phi$ of the $\ast$–FODC leaves $\mathfrak{qg}^{\#}$ invariant ([6]); so
(10) $\mathrm{ad}:={{}_{\Gamma}}\Phi|_{\mathfrak{qg}^{\#}}:\mathfrak{qg}^{\#}\longrightarrow\mathfrak{qg}^{\#}\otimes H$
is a $\mathcal{G}$–corepresentation and it fulfills [6]
(11) $\mathrm{ad}\circ\pi=(\pi\otimes\mathrm{id}_{H})\circ\mathrm{Ad}.$
On the other hand there is a right $H$–module structure on $\mathfrak{qg}^{\#}$ given by
(12) $\theta\circ g=S(g^{(1)})\theta g^{(2)}=\pi(hg-\epsilon(h)g)$
for every $\theta=\pi(h)$in $\mathfrak{qg}^{\#}$; and it satisfies [6]
$(\theta\circ g)^{\ast}=\theta^{\ast}\circ S(g)^{\ast}.$
Let $(\Gamma,d)$ be a $\ast$–FODC over $\mathcal{G}$. Consider the space
(13) $\Gamma^{\wedge}:=\otimes^{\bullet}_{H}\Gamma/\mathcal{Q}\,,\quad\otimes^{\bullet}_{H}\Gamma:=\bigoplus_{k}(\otimes^{k}_{H}\Gamma)\quad\mbox{ with }\quad\otimes^{k}_{H}\Gamma:=\underbrace{\Gamma\otimes_{H}\cdots\otimes_{H}\Gamma}_{k\;times},$
where $\mathcal{Q}$ is the bilateral ideal of $\otimes^{\bullet}_{H}\Gamma$ generated by
(14) $\sum_{i}dg_{i}\otimes_{H}dh_{i}\quad\mbox{ such that }\quad\sum_{i}g_{i}\,dh_{i}=0,$
for $g_{i}$, $h_{i}$in $H$. On the other hand, for a given $t=\vartheta_{1}\cdots\vartheta_{n}$in $\Gamma^{\wedge\,n}$ the linear map
(15) $d:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}$
[…]
Let $(\Gamma,d)$ be a bicovariant $\ast$–FODC. Then the coproduct can be extended to a graded differential $\ast$–algebra morphism
(16) $\Delta:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}\otimes\Gamma^{\wedge}.$
The counit and the coinverse can also be extended to $(\Gamma^{\wedge},d,\ast)$. These maps define a graded differential $\ast$–Hopf algebra structure
(17) $\Gamma^{\wedge\,\infty}:=(\Gamma^{\wedge},\cdot,\mathbbm{1},\Delta,\epsilon,S,d,\ast),$
on $\Gamma^{\wedge}$ which extends $H^{\infty}=(H,\cdot,\mathbbm{1},\Delta,\epsilon,S,d,\ast)$ [4]. In particular,
(18) $\epsilon:\Gamma^{\wedge}\longrightarrow\mathbb{C}$
given by $\epsilon|_{H}:=\epsilon$, $\epsilon|_{\Gamma^{\wedge k}}:=0$ for $k\geq 1$.
Now one can consider the right adjoint action of $\Gamma^{\wedge}$ by
(19) $\mathrm{Ad}:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}\otimes\Gamma^{\wedge}$
such that $\mathrm{Ad}(\vartheta)=(-1)^{\partial\vartheta^{(1)}\partial\vartheta^{(2)}}\vartheta^{(2)}\otimes S(\vartheta^{(1)})\vartheta^{(3)},$ where $\partial x$ denotes the grade of $x$ and $(\mathrm{id}_{\Gamma^{\wedge}}\otimes\Delta)\Delta(\vartheta)=(\Delta\otimes\mathrm{id}_{\Gamma^{\wedge}})\Delta(\vartheta)=\vartheta^{(1)}\otimes\vartheta^{(2)}\otimes\vartheta^{(3)}.$ Clearly, $\mathrm{Ad}$ extends the right adjoint coaction $\mathrm{Ad}$ of $H$.
Let us define
(20) $\mathfrak{qg}^{\#\wedge}=\otimes^{\bullet}\mathfrak{qg}^{\#}/A^{\wedge},\quad\otimes^{\bullet}\mathfrak{qg}^{\#}:=\oplus_{k}(\otimes^{k}\mathfrak{qg}^{\#})\;\mbox{ with }\;\otimes^{k}\mathfrak{qg}^{\#}:=\underbrace{\mathfrak{qg}^{\#}\otimes\cdots\otimes\mathfrak{qg}^{\#}}_{k\;times},$
where $A^{\wedge}$ is the bilateral ideal of $\otimes^{\bullet}\mathfrak{qg}^{\#}$ generated by
$\pi(g^{(1)})\otimes\pi(g^{(2)})$
with $g$in $\mathcal{R}$ [4]. The space $\mathfrak{qg}^{\#\wedge}$ is a graded differential $\ast$–subalgebra ([4]) and it satisfies $\mathfrak{qg}^{\#\wedge}=\{\theta\in\Gamma^{\wedge}\mid\Phi_{\Gamma^{\wedge}}(\theta)=\mathbbm{1}\otimes\theta\}$. Also one can extend the right $H$–module structure of $\mathfrak{qg}^{\#}$ (see equation (12)) to $\mathfrak{qg}^{\#^{\wedge}}$ by means of
(21) $1\circ g=\epsilon(g),\quad(\theta_{1}\theta_{2})\circ g=(\vartheta_{1}\circ g^{(1)})(\theta_{2}\circ g^{(2)}).$
It is worth mentioning that
(22) $d\pi(g)=-\pi(g^{(1)})\pi(g^{(2)}).$
According to [4] the following identification holds:
(23) $\Gamma^{\wedge}=H\otimes\mathfrak{qg}^{\#\wedge}.$
The following example is one of the purposes of this paper
Example 1.2.
Let $G$ be a compact matrix Lie group ($G\subseteq M_{k}(\mathbb{C}))$), and let $\mathcal{G}$ be its associated quantum group [2]. If $H^{\infty}=(H,\cdot,\mathbbm{1},\Delta,\epsilon,S,\ast,)$ is its dense $\ast$–Hopf algebra, then $H$ is the $\ast$–algebra generated by the smooth $\mathbb{C}$–valued functions
$w_{ij}:G$ $\longrightarrow\mathbb{C}$
$A$ $\longmapsto a_{ij}.$
with $A=(a_{ij})$. The coproduct of an element $g$in $H$ is defined by
$\Delta(g):G\times G\longmapsto\mathbb{C},\qquad\Delta(g)(A,B)=g(AB)\quad\mbox{with }\quad A,\,B\,\in\,G.$
In particular, we have $\Delta(w_{ij})=\displaystyle\sum^{n}_{k=1}w_{ik}\otimes w_{kj}.$ The counit and the coinverse are defined as follows:
$\epsilon:H\longrightarrow\mathbb{C},\qquad g\longmapsto g(e),$
where $e$in $G$ is the identity element; and
$S(g):G\longrightarrow\mathbb{C},\qquad A\longmapsto S(g)(A)=g(A^{-1})$
for all $g$in $H$.
Since $G$ is parallelizable, we have the identification $G\times\mathfrak{g}=TG$, where $\mathfrak{g}$ is the Lie algebra of $G$. This diffeomorphism can be given by
(24) $(A,v)\longmapsto(dL_{A})_{e}(v),$
where $(dL_{A})_{e}$ is the differential at $e$ of the left product on $G$ with $A$; and by
(25) $(A,v)\longmapsto(dR_{A})_{e}(v),$
where $(dR_{A})_{e}$ is the differential at $e$ of the right product on $G$ with $A$.
Under equation (24) we have $dg=g^{(1)}\otimes dg^{(2)}_{e}$ for all $g$in $H$, with $dh_{e}$ the differential of $h$in $H$ at $e$. By defining
$\mathfrak{h}:=\mathrm{span}_{\mathbb{C}}\{dg_{e}\mid g\,\in\,H\}$
we obtain that
$(\Gamma=H\otimes\mathfrak{h},d)$
is a $\ast$–FODC over $\mathcal{G}$, where $d$ is the usual differential of smooth $\mathbb{C}$–valued functions. Moreover, it is left–covariant under the map $\Phi_{\Gamma}=\Delta\otimes\mathrm{id}_{\mathfrak{h}}$.
On the other hand, under equation (25), we have $dg=dg^{(1)}_{e}\otimes g^{(2)}$ for all $g$in $H$. Thus
$(\Gamma,d)=(\mathfrak{h}\otimes H,d)$
and it follows that $(\Gamma,d)$ is right–covariant by the map ${{}_{\Gamma}}\Phi=\mathrm{id}_{\mathfrak{h}}\otimes\Delta.$ Hence, $(\Gamma,d)$ is bicovariant. According to [4, 6], there exists a right $H$–ideal $\mathcal{R}\subseteq\mathrm{Ker}(\epsilon)$ such that
$\{\theta\in\Gamma\mid\Phi_{\Gamma}(\theta)=\mathbbm{1}\otimes\theta\}=\mathbbm{1}\otimes{\mathrm{Ker}(\epsilon)\over\mathcal{R}}={\mathrm{Ker}(\epsilon)\over\mathcal{R}}.$
However, since $\mathfrak{h}\cong\mathbbm{1}\otimes\mathfrak{h}=\{\theta\in\Gamma\mid\Phi_{\Gamma}(\theta)=\mathbbm{1}\otimes\theta\}$ it follows that
$\mathfrak{h}={\mathrm{Ker}(\epsilon)\over\mathcal{R}}=\mathfrak{qg}^{\#}.$
We claim that $\mathcal{R}=\mathrm{Ker}^{2}(\epsilon):=\{\displaystyle\sum^{n}_{i=1}a_{i}\,b_{i}\mid a_{i},\,b_{i}\,\in\,\mathrm{Ker}(\epsilon)\,\mbox{ for some }\,n\,\in\,\mathbb{N}\}$. In fact, consider the quantum germs map
$\pi:\mathrm{Ker}(\epsilon)$ $\longrightarrow{\mathrm{Ker}(\epsilon)\over\mathcal{R}}$
$g$ $\longmapsto[g]_{\mathcal{R}}.$
Then, by equation (9), we obtain $\pi(g)=S(g^{(1)})dg^{(2)}=dg_{e}$. If $\mathfrak{g}^{\#}_{\mathbb{C}}$ is the complexification of the dual space of $\mathfrak{g}$, it is well–known that for $X=\{f\in C^{\infty}_{\mathbb{C}}(G)\mid\epsilon(f)=0\}$ we have
$\mathfrak{g}^{\#}_{\mathbb{C}}={X\over X^{2}}$
with $X^{2}=\{\displaystyle\sum^{n}_{i=1}a_{i}\,b_{i}\mid a_{i},\,b_{i}\,\in\,X\,\mbox{ for some }\,n\,\in\,\mathbb{N}\}$ [7]. Moreover, for every $f$in $X$,
$df_{e}=[f]_{X^{2}},$
where $[f]_{X^{2}}$ denotes the equivalence class of $f$ in $X/X^{2}$ [7]. In this way, for every $g$in $\mathrm{Ker}(\epsilon)$ we get $dg_{e}=[g]_{\mathcal{R}}=[g]_{X^{2}}$. By this relation, $g$in $\mathcal{R}$ if and only if $g$in $=X^{2}\cap\mathrm{Ker}(\epsilon)=\mathrm{Ker}^{2}(\epsilon),$ which proves that $\mathcal{R}=\mathrm{Ker}^{2}(\epsilon)$.
By definition, $\mathfrak{h}\subseteq\mathfrak{g}_{\mathbb{C}}$. Since $\mathfrak{g}_{\mathbb{C}}\subseteq M_{k}(\mathbb{C})$, there exists a linear basis $\{v_{i}\}$ of $\mathfrak{g}_{\mathbb{C}}$ composed of linear combinations of the canonical linear basis $\{E_{ij}\}$ of $M_{k}(\mathbb{C})$ ($E_{ij}$ is the matrix with $1$ in the $(i,j)$ position and 00 elsewhere). Then $(dw_{ij})_{e}(v_{k})$ corresponds to the $(i,j)$ position of the matrix $v_{k}$. Consequently, there exist elements $\{(dg_{i})_{e}\}$, which are linear combinations of $(dw_{ij})_{e}$in $\mathfrak{h}$, such that $(dg_{i})_{e}(v_{j})=\delta_{ij}$ with $\delta_{ij}$ being the Kronecker delta. Thus, $\mathfrak{g}^{\#}_{\mathbb{C}}=\mathfrak{h}=\mathfrak{qg}^{\#}$ and we obtain
$(\Gamma,d)=(H\otimes\mathfrak{g}^{\#}_{\mathbb{C}},d).$
This implies that $(\Gamma,d)$ is a $\ast$–subFODC of the $\ast$–FODC over $C^{\infty}_{\mathbb{C}}(G)$
$(C^{\infty}_{\mathbb{C}}(G)\otimes\mathfrak{g}^{\#}_{\mathbb{C}},d)$
of $\mathbb{C}$–valued differential $1$–forms of $G$. Notice that it suffices to take convergent sequences $\{g_{i}\}^{\infty}_{i=1}$subseteq $H$ in $(\Gamma,d)$ to recover $(C^{\infty}_{\mathbb{C}}(G)\otimes\mathfrak{g}^{\#}_{\mathbb{C}},d)$, although this lies beyond our purely geometric–algebraic approach. It is worth mentioning that for the specific form $\Gamma=H\otimes\mathfrak{g}^{\#}_{\mathbb{C}}$, the right covariance map ${{}_{\Gamma}}\Phi$ is given by [6]
${{}_{\Gamma}}\Phi(g\otimes\pi(h))=g^{(1)}\otimes\pi(h^{(2)})\otimes g^{(2)}S(h^{(1)})h^{(1)}$
and hence the $\mathcal{G}$–corepresentation $\mathrm{ad}$ (see equation (10)) coincides with the dualization of the right adjoint action of $G$ on $\mathfrak{g}_{\mathbb{C}}$.
By the Leibniz rule, we have
$\pi(hg)=\pi(h)\epsilon(g)+\epsilon(h)\pi(g)$
and the right $H$–module structure on $\mathfrak{g}^{\#}_{\mathbb{C}}$, given by (see equation (12)) $\pi(h)\circ g=\pi(hg-\epsilon(h)g)$ simplifies to
$\pi(h)\circ g=\epsilon(g)\pi(h).$
Now, consider the universal differential envelope $\ast$–calculus
$(\Gamma^{\wedge},d,\ast)$
of $(\Gamma,d)$. If $g=a\,b$in $\mathrm{Ker}^{2}(\epsilon)$, then by equation (12), we obtain
$\pi(g^{(1)})\otimes\pi(g^{(2)})=\pi(a^{(1)}b^{(1)})\otimes\pi(a^{(2)}b^{(2)})$ $=$ $\pi(a^{(1)})\circ b^{(1)}\otimes\pi(a^{(2)})\circ b^{(2)}$
$+$ $\epsilon(a^{(1)})\pi(b^{(1)})\otimes\pi(a^{(2)})\circ b^{(2)}$
$+$ $\pi(a^{(1)})\circ b^{(1)}\otimes\epsilon(a^{(2)})\pi(b^{(2)})$
$+$ $\epsilon(a^{(1)})\pi(b^{(1)})\otimes\epsilon(a^{(2)})\pi(b^{(2)})$
$=$ $\pi(b)\otimes\pi(a)+\pi(a)\otimes\pi(b).$
and therefore (see equation (20))
$\mathfrak{g}^{\#\,\wedge}_{\mathbb{C}}=\otimes^{\bullet}\mathfrak{g}^{\#}_{\mathbb{C}}/A^{\wedge}=\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}}.$
Thus (see equation (23))
$(\Gamma^{\wedge}=H\otimes\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}},d,\ast).$
By the axioms of the exterior derivative, we can conclude that $d:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}$ is precisely the exterior derivative. Hence $(\Gamma^{\wedge},d,\ast)$ is a graded differential $\ast$–subalgebra of the graded differential $\ast$–algebra
$(C^{\infty}_{\mathbb{C}}(G)\otimes\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}},d,\ast)$
of $\mathbb{C}$–valued differential forms of $G$. As before, it suffices to consider convergent sequences $\{g_{i}\}^{\infty}_{i=1}$subseteq $H$ in $(\Gamma^{\wedge},d,\ast)$ to obtain $(C^{\infty}_{\mathbb{C}}(G)\otimes\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}},d,\ast),$ although this lies beyond our purely geometric–algebraic approach.
The map $\Delta$ is a graded differential $\ast$–morphism. Thus, we have
$\Delta(g_{0}\,\wedge dg_{1}\wedge\cdots\wedge dg_{n})=\Delta(g_{0})\cdot d\Delta(g_{1})\cdots d\Delta(g_{n}).$
Moreover,
$\epsilon(g_{0}\,\wedge dg_{1}\wedge\cdots\wedge dg_{n})=\epsilon(g_{0})d\epsilon(g_{1})\cdots d\epsilon(g_{n})=0,$
as expected. Since $S$ is a graded antimultiplicative map that commutes with $d$, it follows that
$S(g_{0}\,\wedge dg_{1}\wedge\cdots\wedge dg_{n})=(-1)^{n(n-1)\over 2}dS(g_{n})\wedge\cdots\wedge dS(g_{1})\cdot S(g_{0}).$
2. An example about the space of base forms.
Let $(B,\cdot,\mathbbm{1},\ast)$ be a quantum space and let $\mathcal{G}$ be a quantum group. A quantum principal $\mathcal{G}$–bundle over $B$ (abbreviated ”qpb”) is a quantum structure formally represented by the triplet
(26) $\zeta=(P,B,\Delta_{P}),$
where $(P,\cdot,\mathbbm{1},\ast)$ is called the quantum total space, and $(B,\cdot,\mathbbm{1},\ast)$ a quantum subspace, which receives the name quantum base space. Furthermore,
$\Delta_{P}:P\longrightarrow P\otimes H$
is a $\ast$–algebra morphism that satisfies
  1. (1)
    $\Delta_{P}$ is a $\mathcal{G}$–corepresentation.
  2. (2)
    $\Delta_{P}(x)=x\otimes\mathbbm{1}$ if and only if $x$in $B$.
  3. (3)
    The linear map $\beta:P\otimes P\longrightarrow P\otimes H$ given by
    $\beta(x\otimes y):=x\cdot\Delta_{P}(y):=(x\otimes\mathbbm{1})\cdot\Delta_{P}(y)$
    is surjective.
In general, one does not need a quantum group, a $\ast$–Hopf algebra suffices [5, 6]. Given $\zeta$ a qpb over $B$, a differential calculus on it is:
  1. (1)
    A graded differential $\ast$–algebra $(\Omega^{\bullet}(P),d,\ast)$ generated by $\Omega^{0}(P)=P$ (quantum differential forms of $P$).
  2. (2)
    A bicovariant $\ast$–FODC $(\Gamma,d)$ over $\mathcal{G}$.
  3. (3)
    The map $\Delta_{P}$ is extendible to a graded differential $\ast$–algebra morphism
    $\Delta_{\Omega^{\bullet}(P)}:\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}.$
    Here we have considered that $\otimes$ is the tensor product of graded differential $\ast$–algebras.
Note that if $\Delta_{\Omega^{\bullet}(P)}$ exists, it is unique because our graded differential $\ast$–algebras are generated by its degree 00 elements. Furthermore, $\Delta_{\Omega^{\bullet}(P)}$ is a graded differential $\Gamma^{\wedge}$–corepresentation on $\Omega^{\bullet}(P)$ [5]. In this setting, the space of horizontal forms is defined as
(27) $\mathrm{Hor}^{\bullet}P\,:=\{\varphi\in\Omega^{\bullet}(P)\mid\Delta_{\Omega^{\bullet}(P)}(\varphi)\,\in\,\Omega^{\bullet}(P)\otimes H\},$
and it is a graded $\ast$–subalgebra of $\Omega^{\bullet}(P)$ [6]. Since $\Delta_{\Omega^{\bullet}(P)}(\mathrm{Hor}^{\bullet}P)\subseteq\mathrm{Hor}^{\bullet}P\otimes H,$ the map
(28) $\Delta_{\mathrm{Hor}}:=\Delta_{\Omega^{\bullet}(P)}|_{\mathrm{Hor}^{\bullet}P}:\mathrm{Hor}^{\bullet}P\longrightarrow\mathrm{Hor}^{\bullet}P\otimes H$
is $\mathcal{G}$–corepresentation on $\mathrm{Hor}^{\bullet}P$. Also, one can define the space of base forms (quantum differential forms of $B$) as
(29) $\Omega^{\bullet}(B):=\{\mu\in\Omega^{\bullet}(P)\mid\Delta_{\Omega^{\bullet}(P)}(\mu)=\mu\otimes\mathbbm{1}\}.$
The space of base forms is a graded differential $\ast$–subalgebra of $(\Omega^{\bullet}(P),d,\ast)$, and in general, it is not generated by $B$. In fact, the following example, one of the main points of this paper, illustrates a situation in which $(\Omega^{\bullet}(B),d,\ast)$ is not generated by $B$.
Example 2.1.
Let $\mathcal{G}$ be the quantum group associated with $U(1)$ and let
$(H=\mathbb{C}[z,z^{\ast}],\cdot,\mathbbm{1},\Delta,\epsilon,S,\ast)$
[…]
$\zeta=(P:=H,B,\Delta_{P}:=\Delta),$
which is a quantum principal $\mathcal{G}$–bundle over $B$.
For quantum differential forms of $P$, we will use the universal differential envelope $\ast$–calculus $(\Omega^{\bullet}(P),d,\ast)$ presented in Example 1.2 for $U(1)$. Recall that
${{}_{\mathrm{inv}}}\Omega^{1}(P):={\mathrm{Ker}(\epsilon)\over\mathrm{Ker}^{2}(\epsilon)}=\mathfrak{u}(1)^{\#}_{\mathbb{C}}$
is the complexification of the dual space of $\mathfrak{u}(1)$, the Lie algebra of $U(1)$. Of course in this case, $\Omega^{\bullet}(P)$ has no elements of degree $n\geq 2$, and
${{}_{\mathrm{inv}}}\Omega^{1}(P)=\mathrm{span}_{\mathbb{C}}\{\pi(z)=dz_{e}\}.$
In accordance with [6], the set $\{\pi(z)\}$ is a left $P$–basis of $\Omega^{1}(P)$.
For quantum differential forms of $\mathcal{G}$, we will take the universal differential envelope $\ast$–calculus of the bicovariant $\ast$–FODC over $\mathcal{G}$ defined by
$\mathcal{R}=\mathrm{Ker}(\epsilon).$
[…]
Finally, we define
$\Delta_{\Omega^{\bullet}(P)}:\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)\otimes H$
to coincide with $\Delta_{P}$ on the degree 00 case, and for degree $1$ elements,
$\Delta_{\Omega^{\bullet}(P)}(g\,\pi(z))=\Delta(g)(\pi(z)\otimes\mathbbm{1})=g^{(1)}\,\pi(z)\otimes g^{(2)}$
for all $g$in $P$. These constructions provide a differential calculus on $\zeta$. Consequently,
$\Omega^{\bullet}(B)=B\oplus\Omega^{1}(B)\qquad\mbox{ with }\qquad\Omega^{1}(B)=\{\lambda\,\pi(z)\mid\lambda\in\mathbb{C}\}.$
However, this space is not generated by $B$ because $dB=\{0\}$.
3. The Group Isomorphism
In this section, we will consider
Remark 3.1.
Let $\zeta$ be a qpb. Henceforth, we are going to consider that the quantum base space $(B,\cdot,\mathbbm{1},\ast)$ is a $C^{\ast}$–algebra (or it can be completed to a $C^{\ast}$–algebra). Under this assumption and in accordance with [8], for every $\delta^{V}$in $\mathcal{T}$ there exists
$\{T^{\mathrm{L}}_{k}\}^{d_{V}}_{k=1}\subseteq\textsc{Mor}(\delta^{V},\Delta_{P})$
for some $d_{V}$in $\mathbb{N}$ such that
(30) $\sum^{d_{V}}_{k=1}x^{V\,\ast}_{ki}x^{V}_{kj}=\delta_{ij}\mathbbm{1},$
with $x^{V}_{ki}:=T^{\mathrm{L}}_{k}(e_{i})$, where $\mathcal{T}$ is a complete set of mutually non–equivalent irreducible (necessarily finite–dimensional) $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$ (the trivial corepresentation on $\mathbb{C}$) and $\{e_{i}\}^{n_{V}}_{i=1}$ is the orthonormal basis of $V$ shown in Theorem 1.1.
Let $\zeta=(P,B,\Delta_{P})$ be a qpb. The surjective map $\beta$ can be used to define the linear isomorphism
$\widetilde{\beta}:P\otimes_{B}P\longrightarrow P\otimes H$
such that $\widetilde{\beta}(x\otimes_{B}y)=\beta(x\otimes y)=(x\otimes\mathbbm{1})\cdot\Delta_{P}(y).$ [6]. The degree zero quantum translation map is defined as
(31) $\mathrm{qtrs}:H\longrightarrow P\otimes_{B}P,$
such that $\mathrm{qtrs}(g)=\widetilde{\beta}^{-1}(\mathbbm{1}\otimes g).$ Explicitly, by taking the linear basis $\{g^{V}_{ij}\}_{\delta^{V},i,j}$ (see Theorem 1.1 and Remark 3.1), we have
(32) $\mathrm{qtrs}(g^{V}_{ij})=\sum^{d_{V}}_{k=1}x^{V\,\ast}_{ki}\otimes_{B}x^{V}_{kj}.$
In particular, since $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$ we have that $\mathbbm{1}$in $\{g^{V}_{ij}\}_{\delta^{V},i,j}$ and
$\mathrm{qtrs}(\mathbbm{1})=\mathbbm{1}\otimes_{B}\mathbbm{1}.$
We can extend $\mathrm{qtrs}$ to
(33) $\widetilde{\mathrm{qtrs}}:P\otimes H\longrightarrow P\otimes_{B}P$
by means of $\widetilde{\mathrm{qtrs}}(x\otimes g^{V}_{ij})=x\,\mathrm{qtrs}(g^{V}_{ij})=\displaystyle\sum^{d_{\alpha}}_{k=1}x\,x^{V\,\ast}_{ki}\otimes_{B}x^{V}_{kj}.$ A direct calculation shows that $\widetilde{\beta}$ and $\widetilde{\mathrm{qtrs}}$ are mutually inverse.
Throughout the various computations of this paper, we shall use the symbolic notation
(34) $\mathrm{qtrs}(g)=[g]_{1}\otimes_{B}[g]_{2}.$
Now we shall assume that $\zeta=(P,B,\Delta)$ is endowed with a differential calculus. In this situation $\widetilde{\beta}$ has a natural extension to
(35) $\widetilde{\beta}:\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}$
given by $\widetilde{\beta}(w_{1}\otimes_{\Omega^{\bullet}(B)}w_{2})=(w_{1}\otimes\mathbbm{1})\cdot\Delta_{\Omega^{\bullet}(P)}(w_{2}),$ where the tensor product on the image is the tensor product of graded differential $\ast$–algebras. According to [10], this map is bijective.
On the other hand, taking a real qpc $\omega$ (which always exists [5]) and in accordance with [10], we can extend $\mathrm{qtrs}$ to
(36) $\mathrm{qtrs}:\Gamma\longrightarrow\left(\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P)\right)^{1}$
by means of
(37) $\mathrm{qtrs}(\theta)=\mathbbm{1}\otimes_{\Omega^{\bullet}(B)}\omega(\theta)-(m_{\Omega}\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{P})(\omega\otimes\mathrm{qtrs})\mathrm{ad}(\theta)$
when $\theta$in $\mathfrak{qg}^{\#}$, where
$m_{\Omega}:\Omega^{\bullet}(P)\otimes\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)$
is the product map; and
(38) $\mathrm{qtrs}(\vartheta\upsilon):=(-1)^{\partial\vartheta\,\partial[\upsilon]_{1}}[\upsilon]_{1}\,\mathrm{qtrs}(\vartheta)\,[\upsilon]_{2}$
for $\vartheta$in $H$, $\upsilon$in $\mathfrak{qg}^{\#}$ or $\vartheta$in $\mathfrak{qg}^{\#}$, $\upsilon$in $H$, with $\mathrm{qtrs}(\upsilon)=[\upsilon]_{1}\otimes_{\Omega^{\bullet}(B)}[\upsilon]_{2}$, where $\partial\vartheta$ is the grade of $\vartheta$. A direct calculation shows that the corresponding extended map $\widetilde{\mathrm{qtrs}}$ is the right inverse of $\widetilde{\beta}$ for degree $1$ and since $\widetilde{\beta}$ is bijective, we get that $\widetilde{\mathrm{qtrs}}$ is actually its inverse. It is worth mentioning that even when apparently the definition of $\mathrm{qtrs}$ depends on the real qpc $\omega$ chosen, by the uniqueness of the inverse function, the last result tells us that $\mathrm{qtrs}$ is independent of this choice. Since $\widetilde{\beta}$ commutes with the corresponding differential maps, it follows that ([10])
Proposition 3.2.
We have
$\mathrm{qtrs}\circ d=d_{\otimes^{\bullet}}\circ\mathrm{qtrs},$
where $d_{\otimes^{\bullet}}$ is the differential map of $\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P)$.
Now let us take the universal graded differential $\ast$ –algebra $(\otimes^{\bullet}_{H}\Gamma,d_{\otimes_{H}},\ast)$ of $(\Gamma,d)$ [6]. The quantum translation map can be extended naturally to $\otimes^{\bullet}_{H}\Gamma$ by means of
(39) $\mathrm{qtrs}(\vartheta\otimes_{H}\upsilon):=(-1)^{\partial\vartheta\,\partial[\upsilon]_{1}}\,[\upsilon]_{1}\,\mathrm{qtrs}(\vartheta)\,[\upsilon]_{2}$
if $\mathrm{qtrs}(\upsilon)=[\upsilon]_{1}\otimes_{\Omega^{\bullet}(B)}[\upsilon]_{2}$. By induction and Proposition 3.2 it can be proved that
(40) $\mathrm{qtrs}(dg_{0}\otimes_{H}dg_{1}\otimes_{H}...\otimes_{H}dg_{k})=d_{\otimes^{\bullet}}(\mathrm{qtrs}(g_{0}\,dg_{1}\otimes_{H}...\otimes_{H}dg_{k}))$
for $g_{1}$,…, $g_{k}$in $H$. Now by considering the bilateral ideal $\mathcal{Q}$ generated by the relation presented on equation (14), a direct calculation shows that $\mathrm{qtrs}(\mathcal{Q})=0$ and therefore, $\mathrm{qtrs}$ can be defined on the universal differential envelope $\ast$–calculus
(41) $\mathrm{qtrs}:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P).$
As before, the corresponding extended map
(42) $\widetilde{\mathrm{qtrs}}:\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P).$
given by $\widetilde{\mathrm{qtrs}}(w_{1}\otimes\vartheta)=(w_{1}\otimes\mathbbm{1})\cdot\mathrm{qtrs}(\vartheta)$ is the inverse map of $\widetilde{\beta}$ [10]. According to [11] we have
Proposition 3.3.
The following properties hold
  1. (1)
    $[\vartheta]_{1}\,[\vartheta]_{2}=\epsilon(\vartheta)\mathbbm{1}$.
  2. (2)
    $(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\Delta_{\Omega^{\bullet}(P)})\circ\mathrm{qtrs}=(\mathrm{qtrs}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\Delta$.
  3. (3)
    $(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{\Omega^{\bullet}(P)})\circ\mathrm{qtrs}=(\sigma\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{\Omega^{\bullet}(P)})\circ(S\otimes\mathrm{qtrs})\circ\Delta$, where
    $\sigma:\Gamma^{\wedge}\otimes\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}$
    is the canonical graded twist map, i.e., $\sigma(\vartheta\otimes w)=(-1)^{kl}\,w\otimes\vartheta$ if $w$in $\Omega^{k}(P)$ and $\vartheta$in $\Gamma^{\wedge l}$.
  4. (4)
    $\mu\,\mathrm{qtrs}(\vartheta)=(-1)^{lk}\mathrm{qtrs}(\vartheta)\,\mu$ for all $\mu$in $\Omega^{k}(B)$, $\vartheta$in $\Gamma^{\wedge l}.$
Let
$\mathfrak{f}_{1},\,\mathfrak{f}_{2}:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P)$
be two graded linear maps. The convolution product of $\mathfrak{f}_{1}$ with $\mathfrak{f}_{2}$ is defined by
$\mathfrak{f}_{1}\ast\mathfrak{f}_{2}=m_{\Omega}\circ(\mathfrak{f}_{1}\otimes\mathfrak{f}_{2})\circ\Delta:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P).$
Henceforth we will just consider graded maps $\mathfrak{f}$ such that
(43) $\mathfrak{f}(\mathbbm{1})=\mathbbm{1}\qquad\mbox{and}\qquad(\mathfrak{f}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\mathrm{Ad}=\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{f},$
where $\mathrm{Ad}:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}\otimes\Gamma^{\wedge}$ is the extension of the right $\mathcal{G}$–coaction $\mathrm{Ad}:H\longrightarrow H\otimes H$ (see equation (19)). We say that $\mathfrak{f}$ is a convolution invertible map if there exists a graded linear map
$\mathfrak{f}^{-1}:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P)$
such that
(44) $\mathfrak{f}\ast\mathfrak{f}^{-1}=\mathfrak{f}^{-1}\ast\mathfrak{f}=\mathbbm{1}\epsilon.$
A direct calculation shows that the set of all convolution invertible maps $\{\mathfrak{f}:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P)\}$ is a group with respect to the convolution product. The next proposition is one of the purposes of this paper
Proposition 3.4.
There exist a group isomorphism between the group of all convolution invertible maps and the group of all graded left $\Omega^{\bullet}(B)$–module isomorphisms
$\mathfrak{F}:\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)$
that satisfy
(45) $\mathfrak{F}(\mathbbm{1})=\mathbbm{1}\qquad\mbox{ and }\qquad(\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\Delta_{\Omega^{\bullet}(P)}=\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{F}.$
Here we are considering the group product $(\mathfrak{F}_{1}\circ\mathfrak{F}_{2})(w)=\mathfrak{F}_{2}(\mathfrak{F}_{1}(w))$.
Proof.
Let us start by considering a map $\mathfrak{F}$. Then we define a graded linear map
(46) $\mathfrak{f}_{\mathfrak{F}}:=m_{\Omega^{\bullet}}\circ(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F})\circ\mathrm{qtrs}:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P),$
where $m_{\Omega^{\bullet}}:\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)$ is the product map. We are going to show that $\mathfrak{f}_{\mathfrak{F}}$ is a convolution invertible map. First, since $\mathrm{qtrs}(\mathbbm{1})=\mathbbm{1}\otimes_{B}\mathbbm{1}$ it follows $\mathfrak{f}_{\mathfrak{F}}(\mathbbm{1})=\mathbbm{1}.$ Secondly, according to Proposition 3.3 point $3$
$\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{f}_{\mathfrak{F}}$ $=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}(\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{F}))\circ\mathrm{qtrs}$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\mathrm{id}_{\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}}\otimes_{\Omega^{\bullet}(B)}(\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{F}))\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{\Omega^{\bullet}(P)})\circ\mathrm{qtrs}$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\mathrm{id}_{\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}}\otimes_{\Omega^{\bullet}(B)}(\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{F}))\circ(\sigma\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{\Omega^{\bullet}(P)})\circ(S\otimes\mathrm{qtrs})\circ\Delta,$
where
$\widehat{m}_{\Omega^{\bullet}}:(\Omega^{\bullet}(P)\otimes\Gamma^{\wedge})\otimes_{\Omega^{\bullet}(B)}(\Omega^{\bullet}(P)\otimes\Gamma^{\wedge})\longrightarrow\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}$
is such that $\widehat{m}_{\Omega^{\bullet}}(w_{1}\otimes\vartheta_{1}\otimes_{\Omega^{\bullet}(B)}w_{2}\otimes\vartheta_{2})=(-1)^{kl}w_{1}w_{2}\otimes\vartheta_{1}\vartheta_{2}$ if $w_{2}$in $\Omega^{k}(P)$ and $\vartheta_{1}$in $\Gamma^{\wedge l}$. On the other hand, by equation (45) and Proposition 3.3 point $2$
$\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{f}_{\mathfrak{F}}$ $=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}(\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{F}))\circ\mathrm{qtrs}$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}((\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\Delta_{\Omega^{\bullet}(P)}))\circ\mathrm{qtrs}$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}(\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}}))\circ(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\Delta_{\Omega^{\bullet}(P)})\circ\mathrm{qtrs}$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}(\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}}))\circ(\mathrm{qtrs}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\Delta;$
but by considering again Proposition 3.3 point $3$
$\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{f}_{\mathfrak{F}}$ $=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}(\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}}))\circ(\mathrm{qtrs}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\Delta$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\mathrm{id}_{\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}}\otimes_{\Omega^{\bullet}(B)}(\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}}))\circ(\Delta_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}})$
$\circ\,(\mathrm{qtrs}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\Delta$
$=$ $\widehat{m}_{\Omega^{\bullet}}\circ(\mathrm{id}_{\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}}\otimes_{\Omega^{\bullet}(B)}(\mathfrak{F}\otimes\mathrm{id}_{\Gamma^{\wedge}}))$
$\circ\,[(({\sigma}\otimes_{\Omega^{\bullet}(B)}\mathrm{id}_{\Omega^{\bullet}(P)})\circ(S\otimes\mathrm{qtrs})\circ\Delta)\otimes\mathrm{id}_{\Gamma^{\wedge}}]\circ\Delta.$
So for all $\vartheta$in $\Gamma^{\wedge}$
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{f}_{\mathfrak{F}}(\vartheta))$ $=$ $(-1)^{\partial\vartheta^{(1)}(\partial[\vartheta^{(2)}]_{1}+\partial[\vartheta^{(2)}]_{2})}\,[\vartheta^{(2)}]_{1}F([\vartheta^{(2)}]_{2})\otimes S(\vartheta^{(1)})\vartheta^{(3)}$
$=$ $(-1)^{\partial\vartheta^{(1)}\partial\vartheta^{(2)}}\,[\vartheta^{(2)}]_{1}F([\vartheta^{(2)}]_{2})\otimes S(\vartheta^{(1)})\vartheta^{(3)}\;=\;(\mathfrak{f}_{\mathfrak{F}}\otimes\mathrm{id}_{\Gamma^{\wedge}})\mathrm{Ad}(\vartheta),$
because $\mathrm{qtrs}$ is a graded linear map. Finally, consider
$\mathfrak{f}_{\mathfrak{F}^{-1}}:=m_{\Omega^{\bullet}}\circ(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\circ\mathrm{qtrs}:\Gamma^{\wedge}\longrightarrow\Omega^{\bullet}(P).$
Then for all $\vartheta$in $\Gamma^{\wedge}$
$(\mathfrak{f}_{\mathfrak{F}}\ast\mathfrak{f}_{\mathfrak{F}^{-1}})(\vartheta)=[\vartheta^{(1)}]_{1}\,\underbrace{\mathfrak{F}([\vartheta^{(1)}]_{2})\,[\vartheta^{(2)}]_{1}}\,\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2}).$
We claim that the expression in the brace is an element of $\Omega^{\bullet}(B)$. Indeed, notice that
(49) $\Delta_{\Omega^{\bullet}(P)}((\mathfrak{f}_{\mathfrak{F}}\ast\mathfrak{f}_{\mathfrak{F}^{-1}})(\vartheta))$ $=$ $\Delta_{\Omega^{\bullet}(P)}\left([\vartheta^{(1)}]_{1}\,\mathfrak{F}([\vartheta^{(1)}]_{2})\,[\vartheta^{(2)}]_{1}\,\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2})\right)$
$=$ $\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([\vartheta^{(1)}]_{2})\,[\vartheta^{(2)}]_{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2})).$
Also we know
$\Delta_{\Omega^{\bullet}(P)}((\mathfrak{f}_{\mathfrak{F}}\ast\mathfrak{f}_{\mathfrak{F}^{-1}})(\vartheta))=\Delta_{\Omega^{\bullet}(P)}(\mathfrak{f}_{\mathfrak{F}}(\vartheta^{(1)}))\cdot\Delta_{\Omega^{\bullet}(P)}(\mathfrak{f}_{\mathfrak{F}^{-1}}(\vartheta^{(2)})).$
By applying equation (3) on $\Delta_{\Omega^{\bullet}(P)}(\mathfrak{f}_{\mathfrak{F}}(\vartheta^{(1)}))$ and equation (3) on $\Delta_{\Omega^{\bullet}(P)}(\mathfrak{f}_{\mathfrak{F}}(\vartheta^{(2)}))$ we have
$\Delta_{\Omega^{\bullet}(P)}((\mathfrak{f}_{\mathfrak{F}}\ast\mathfrak{f}_{\mathfrak{F}^{-1}})(\vartheta))$ $=$ $(-1)^{\partial\vartheta^{(3)}\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\vartheta^{(2)})$
$([\vartheta^{(4)}]_{1}\otimes S(\vartheta^{(3)}))\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(4)}]_{2}))$
$=$ $(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(4)}]_{1}\otimes\vartheta^{(2)}S(\vartheta^{(3)}))$
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(4)}]_{2})).$
On the other hand, by the definition of $\epsilon$ (see equation (18)) and the graded Hopf $\ast$–algebra structure of $\Gamma^{\wedge\,\infty}$ and we get
$\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(2)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2}))$
$=\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\mathbbm{1})\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(2)}))$
[…]
$=(-1)^{\partial\vartheta^{(2)}\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\;\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\underbrace{\epsilon(\vartheta^{(2)})\mathbbm{1}}_{0-\mathrm{degree}})\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(3)})).$
$=(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\underbrace{\vartheta^{(2)}S(\vartheta^{(3)}}_{0-\mathrm{degree}}))\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F}^{-1})\mathrm{qtrs}(\vartheta^{(4)}))$
$=(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})\otimes\underbrace{\vartheta^{(2)}S(\vartheta^{(3)}}_{0-\mathrm{degree}}))\,([\vartheta^{(4)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(4)}]_{2}))$
$=(-1)^{(\partial\vartheta^{(3)}+\partial\vartheta^{(2)})\partial[\vartheta^{(4)}]_{1}}\,\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(4)}]_{1}\otimes\vartheta^{(2)}S(\vartheta^{(3)}))\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(4)}]_{2}))$
and hence
$\Delta_{\Omega^{\bullet}(P)}((\mathfrak{f}_{\mathfrak{F}}\ast\mathfrak{f}_{\mathfrak{F}^{-1}})(\vartheta))=\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(2)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2})).$
By equation (49) and the last equality we conclude
$\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([\vartheta^{(1)}]_{2})\,[\vartheta^{(2)}]_{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2}))$
$=\Delta_{\Omega^{\bullet}(P)}([\vartheta^{(1)}]_{1})\,(\mathfrak{F}([\vartheta^{(1)}]_{2})[\vartheta^{(2)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2})),$
which proves our claim. Since $\mathfrak{F}^{-1}$ is a left $\Omega^{\bullet}(B)$–module morphism, by our previous claim and Proposition 3.3 point $1$
$(\mathfrak{f}_{\mathfrak{F}}\ast\mathfrak{f}_{\mathfrak{F}^{-1}})(\vartheta)\;=\;[\vartheta^{(1)}]_{1}\,\mathfrak{F}([\vartheta^{(1)}]_{2})\,[\vartheta^{(2)}]_{1}\,\mathfrak{F}^{-1}([\vartheta^{(2)}]_{2})$ $=$ $[\vartheta^{(1)}]_{1}\,\mathfrak{F}^{-1}(\mathfrak{F}([\vartheta^{(1)}]_{2})\,[\vartheta^{(2)}]_{1}[\vartheta^{(2)}]_{2})$
$=$ $[\vartheta^{(1)}]_{1}\,\mathfrak{F}^{-1}(\mathfrak{F}([\vartheta^{(1)}]_{2})\,\epsilon(\vartheta^{(2)}))$
$=$ $\epsilon(\vartheta^{(1)})\epsilon(\vartheta^{(2)})\mathbbm{1}$
$=$ $\epsilon(\vartheta)\mathbbm{1}$
for every $\vartheta$in $\Gamma^{\wedge}$. In a similar way, we can prove that $\mathfrak{f}_{\mathfrak{F}^{-1}}\ast\mathfrak{f}_{\mathfrak{F}}=\mathbbm{1}\epsilon$ and hence $\mathfrak{f}_{\mathfrak{F}}$ is a convolution invertible map.
Conversely, for a given convolution invertible map $\mathfrak{f}$ let us define the graded linear map
(50) $\mathfrak{F}_{\mathfrak{f}}:=m_{\Omega}\circ(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})\circ\Delta_{\Omega^{\bullet}(P)}:\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P).$
We are going to prove that $\mathfrak{F}_{\mathfrak{f}}$ is a graded left $\Omega^{\bullet}(B)$–module isomorphism which satisfies equation (45). First of all, by equation (43) it is obvious that $\mathfrak{F}_{\mathfrak{f}}(\mathbbm{1})=\mathbbm{1}$. Secondly, taking $\mu$in $\Omega^{\bullet}(B)$ and $w$in $\Omega^{\bullet}(P)$ we have $\mathfrak{F}_{\mathfrak{f}}(\mu w)=m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})(\mu\otimes\mathbbm{1})\Delta_{\Omega^{\bullet}(P)}(w)=\mu\,\mathfrak{F}_{f}(w).$ Thirdly, by equation (43)
$\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{F}_{\mathfrak{f}}$ $=$ $\widehat{m}_{\Omega}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes(\Delta_{\Omega^{\bullet}(P)}\circ\mathfrak{f}))\circ\Delta_{\Omega^{\bullet}(P)}$
$=$ $\widehat{m}_{\Omega}\circ(\Delta_{\Omega^{\bullet}(P)}\otimes((\mathfrak{f}\otimes\mathrm{id}_{\Gamma^{\wedge}})\circ\mathrm{Ad}))\circ\Delta_{\Omega^{\bullet}(P)},$
where
$\widehat{m}_{\Omega}:(\Omega^{\bullet}(P)\otimes\Gamma^{\wedge})\otimes(\Omega^{\bullet}(P)\otimes\Gamma^{\wedge})\longrightarrow\Omega^{\bullet}(P)\otimes\Gamma^{\wedge}$
[…]
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}_{\mathfrak{f}}(w))$ $=$ $(-1)^{\partial w^{3}(\partial w^{1}+\partial w^{2})}w^{(0)}\mathfrak{f}(w^{(3)})\otimes w^{(1)}S(w^{(2)})w^{(4)}$
$=$ $(-1)^{\partial w^{3}\partial w^{1}}w^{(0)}\mathfrak{f}(w^{(2)})\otimes\epsilon(w^{(1)})w^{(3)}$
$=$ $w^{(0)}\mathfrak{f}(w^{(2)})\otimes\epsilon(w^{(1)})w^{(3)}$
$=$ $w^{(0)}\mathfrak{f}(w^{(2)})\otimes\epsilon(w^{(1)})w^{(3)}$
$=$ $w^{(0)}\mathfrak{f}(w^{(1)})\otimes w^{(2)}=(\mathfrak{F}_{\mathfrak{f}}\otimes\mathrm{id}_{\Gamma^{\wedge}})\Delta_{\Omega^{\bullet}(P)}(w).$
Finally, notice that for all $w$in $\Omega^{\bullet}(P)$
$\mathfrak{F}_{\mathfrak{f}^{-1}}(\mathfrak{F}_{\mathfrak{f}}(w))\;=\;m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f}^{-1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}_{\mathfrak{f}}(w))$ $=$ $m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f}^{-1})(\mathfrak{F}_{\mathfrak{f}}\otimes\mathrm{id}_{\Gamma^{\wedge}})\,\Delta_{\Omega^{\bullet}(P)}(w)$
$=$ $\mathfrak{F}_{f}(w^{(0)})\,\mathfrak{f}^{-1}(w^{(1)})$
$=$ $w^{(0)}\,\mathfrak{f}(w^{(1)})\,\mathfrak{f}^{-1}(w^{(2)})$
$=$ $w^{(0)}\,\epsilon(w^{(1)})\;=\;w.$
A similar calculation shows $\mathfrak{F}_{\mathfrak{f}}(\mathfrak{F}_{\mathfrak{f}^{-1}}(w))=w$, so $\mathfrak{F}_{\mathfrak{f}}$ is biyective and $\mathfrak{F}^{-1}_{\mathfrak{f}}=\mathfrak{F}_{\mathfrak{f}^{-1}}.$
Our next step is to prove that
$\mathfrak{F}\xmapsto{\widehat{\Lambda}}\mathfrak{f}_{\mathfrak{F}}\,,\qquad\mathfrak{f}\xmapsto{\widetilde{\Lambda}}\mathfrak{F}_{\mathfrak{f}}$
are mutually inverse. Notice that for all $w$in $\Omega^{\bullet}(P)$
$\mathfrak{F}_{\mathfrak{f}_{\mathfrak{F}}}(w)=m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes(m_{\Omega^{\bullet}}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F})\mathrm{qtrs}))\Delta_{\Omega^{\bullet}(P)}(w)=\underbrace{w^{(0)}[w^{(1)}]_{1}}\mathfrak{F}([w^{(1)}]_{2}),$
where the expression in the brace is an element of $\Omega^{\bullet}(B)$. In fact,
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}_{\mathfrak{f}_{\mathfrak{F}}}(w))=\Delta_{\Omega^{\bullet}(P)}\left(w^{(0)}[w^{(1)}]_{1}\mathfrak{F}([w^{(1)}]_{2})\right)=\Delta_{\Omega^{\bullet}(P)}(w^{(0)}[w^{(1)}]_{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(1)}]_{2}));$
however by equation (3) we get
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}_{\mathfrak{f}_{\mathfrak{F}}}(w))\;=\;\Delta_{\Omega^{\bullet}(P)}(w^{(0)}\mathfrak{f}_{\mathfrak{F}}(w^{(1)}))$ $=$ $(w^{(0)}\otimes w^{(1)})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{f}_{\mathfrak{F}}(w^{(2)}))$
$=$ $(-1)^{\partial w^{(2)}\partial[w^{(3)}]_{1}}(w^{(0)}\otimes w^{(1)})$
$([w^{(3)}]_{1}\otimes S(w^{(2)}))\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(3)}]_{2}))$
$=$ $(-1)^{(\partial w^{(1)}+\partial w^{(2)})\partial[w^{(3)}]_{1}}\,(w^{(0)}[w^{(3)}]_{1}\otimes w^{(1)}S(w^{(2)}))$
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(3)}]_{2}))$
[…]
$\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(2)}]_{2}))$
$=$ $(w^{(0)}[w^{(2)}]_{1}\otimes\epsilon(w^{(1)})\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(2)}]_{2}))$
$=$ $(w^{(0)}\otimes\epsilon(w^{(1)})\mathbbm{1})([w^{(2)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(2)}]_{2}))$
$=$ $(w^{(0)}\otimes\epsilon(w^{(1)})\mathbbm{1})\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F})\mathrm{qtrs}(w^{(2)}))$
$=$ $(w^{(0)}\otimes\mathbbm{1})\,\widetilde{\beta}((\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}\mathfrak{F})\mathrm{qtrs}(w^{(1)}))$
$=$ $(w^{(0)}[w^{(1)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(1)}]_{2})).$
This implies that
$\Delta_{\Omega^{\bullet}(P)}(w^{(0)}[w^{(1)}]_{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(1)}]_{2}))=(w^{(0)}[w^{(1)}]_{1}\otimes\mathbbm{1})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}([w^{(1)}]_{2})),$
which proves our assertion. Since $\mathfrak{F}$ is a left $\Omega^{\bullet}(B)$–module morphism, by this assertion and Proposition 3.3 point $1$, we have
$\mathfrak{F}_{\mathfrak{f}_{\mathfrak{F}}}(w)=w^{(0)}[w^{(1)}]_{1}\,\mathfrak{F}([w^{(1)}]_{2})=\mathfrak{F}(w^{(0)}[w^{(1)}]_{1}[w^{(1)}]_{2}))=\mathfrak{F}(w^{(0)}\epsilon(w^{(1)}))=\mathfrak{F}(w).$
On the other hand, for every $\vartheta$in $\Gamma^{\wedge}$
$\mathfrak{f}_{\mathfrak{F}_{\mathfrak{f}}}(\vartheta)$ $=$ $m_{\Omega^{\bullet}}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes_{\Omega^{\bullet}(B)}(m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})\Delta_{\Omega^{\bullet}(P)}))\mathrm{qtrs}(\vartheta)$
$=$ $[\vartheta]_{1}[\vartheta]^{(0)}_{2}\mathfrak{f}([\vartheta]^{(1)}_{2})$
$=$ $m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})([\vartheta]_{1}\otimes\mathbbm{1})([\vartheta]^{(0)}_{2}\otimes[\vartheta]^{(1)}_{2})$
$=$ $m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})([\vartheta]_{1}\otimes\mathbbm{1})\Delta_{\Omega^{\bullet}(P)}([\vartheta]_{2})$
$=$ $m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})\,\widetilde{\beta}(\mathrm{qtrs}(\vartheta))$
$=$ $m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f})(\mathbbm{1}\otimes\vartheta)\;=\;\mathfrak{f}(\vartheta)$
and hence $\widetilde{\Lambda}=\widehat{\Lambda}^{-1}$. Finally, to complete the proof it is enough to prove that $\widehat{\Lambda}$ or $\widehat{\Lambda}^{-1}$ is a group morphism. In fact
$\mathfrak{F}_{\mathfrak{f}_{1}\ast\mathfrak{f}_{2}}(w)=m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes(\mathfrak{f}_{1}\ast\mathfrak{f}_{2}))\Delta_{\Omega^{\bullet}(P)}(w)=w^{(0)}\,\mathfrak{f}_{1}(w^{(1)})\,\mathfrak{f}_{2}(w^{(2)});$
while by equation (45)
$\mathfrak{F}_{\mathfrak{f}_{2}}(\mathfrak{F}_{\mathfrak{f}_{1}}(w))\;=\;m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f}_{2})\,\Delta_{\Omega^{\bullet}(P)}(\mathfrak{F}_{\mathfrak{f}_{1}}(w))$ $=$ $m_{\Omega}(\mathrm{id}_{\Omega^{\bullet}(P)}\otimes\mathfrak{f}_{2})(\mathfrak{F}_{\mathfrak{f}_{1}}\otimes\mathrm{id}_{\Gamma^{\wedge}})\,\Delta_{\Omega^{\bullet}(P)}(w)$
$=$ $\mathfrak{F}_{\mathfrak{f}_{1}}(w^{(0)})\,\mathfrak{f}_{2}(w^{(1)})$
$=$ $w^{(0)}\,\mathfrak{f}_{1}(w^{(1)})\,\mathfrak{f}_{2}(w^{(2)})$
for all $w$in $\Omega^{\bullet}(P)$. Therefore $\mathfrak{f}_{1}\ast\mathfrak{f}_{2}\xmapsto{\widehat{\Lambda}^{-1}}\mathfrak{F}_{\mathfrak{f}_{1}}\circ\mathfrak{F}_{\mathfrak{f}_{2}}$ and the proposition follows. ∎
4. The form of the maps $\{T^{\mathrm{L}}_{k}\}$
Let $G$ be a compact matrix Lie group, and let $\mathcal{G}$ be its associated quantum group [2]. If $\delta^{V}$ is a (unitary) irreducible $\mathcal{G}$–corepresentation, then it induces a $G$–representation on $V$. Indeed, by Theorem 1.1 we know that
$\delta^{V}(e_{j})=\sum^{n}_{i=1}e_{i}\otimes g^{V}_{ij}.$
Then the linear map
$\alpha^{V}:G\times V\longrightarrow V$
given by
$\alpha^{V}(A,e_{j})=\sum^{n}_{i=1}g^{V}_{ij}(A)\,e_{i}$
is a unitary and irreducible $G$–representation. The following proposition is another of our purposes.
Proposition 4.1.
Let $G$ be a compact matrix Lie group and let $\pi:P\longrightarrow B$ be a classical principal $G$–bundle, where $P$ is the total space, $B$ is the base space and $\pi$ is the bundle projection. Assume $P$, $B$ are compact. If $\mathcal{T}$ is a complete set of mutually non–equivalent irreducible $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$, then for every $\delta^{V}$in $\mathcal{T}$, there exists $\{T^{\mathrm{L}}_{k}\}^{d_{V}}_{k=1}$subseteq $\textsc{Mor}(\delta^{V},\Delta_{P})$, for some $d_{V}$in $\mathbb{N}$, such that equation (30) holds. Here
$\Delta_{P}:C^{\infty}_{\mathbb{C}}(P)\longrightarrow C^{\infty}_{\mathbb{C}}(P\times G)\cong C^{\infty}_{\mathbb{C}}(P)\otimes C^{\infty}_{\mathbb{C}}(G)$
is the pull–back of the right $G$–action on $P$. In this setting, $C^{\infty}_{\mathbb{C}}(P)$ denotes the space of $\mathbb{C}$–valued smooth functions of $P$ and the tensor product is taken to be the completed injective tensor product.
Proof.
Let $\delta^{V}$in $\mathcal{T}$ and consider its associated representation $\alpha^{V}$ with $n=\mathrm{dim}_{\mathbb{C}}(V)$. Since $\alpha^{V}$ is unitary, the associated vector bundle $\pi_{\alpha^{V}}:E^{V}\longrightarrow B$ is a Hermitian bundle.
For each $b$in $B$, let $(U_{b},\Phi_{b})$ be a local trivialization of the Hermitian bundle $\pi_{\alpha^{V}}:E^{V}\longrightarrow B$ around $b$ associated to a principal $G$–bundle local trivialization $(U_{b},\Psi_{b})$ of $\pi:P\longrightarrow B$. Then there exists a set of local sections $\{\hat{s}^{b}_{1},\ldots,\hat{s}^{b}_{1}\}$subseteq $\Gamma\bigl{(}\pi_{{\alpha^{V}}}^{-1}(U_{b})\bigr{)}$ such that $\hat{s}^{b}_{i}(a)=(a,e_{i})$ for all $a$in $U_{b}$ (under the diffeomorphism $\Phi_{b}$). Since $\{U_{b}\}_{b\in B}$ is an open cover, by compactness, there exist points $b_{1},\ldots,b_{r}\in B$ such that $\{\,U_{b_{i}}\}_{i=1}^{r}$ remains an open cover of $B$. Let $\{\rho_{b_{i}}\}_{i=1}^{r}$ be a partition of unity subordinate to the open cover $\{\,U_{b_{i}}\}_{i=1}^{r}$, where each $\rho_{b_{i}}$ has compact support and admits a smooth square root. In this way, consider the global sections of $E^{V}$
$\{s^{i}_{j}=\sqrt{\rho_{b_{i}}}\,\hat{s}^{b_{i}}_{j}\}^{r,n}_{i,j=1}.$
Notice that $\{s^{i}_{j}\}^{r,n}_{i,j=1}$ is a set of $\mathbb{C}^{\infty}_{\mathbb{C}}(B)$–bimodule generators of $\Gamma(E^{V})$.
It is well–known that $\Gamma(E^{V})$ is isomorphic to the space of $G$–equivariant smooth functions $C^{\infty}_{\mathbb{C}}(P,V)^{G}=\{f:P\longrightarrow V\mid f\mbox{ is smooth such that }f(xA)=\alpha(A^{-1})f(x)\}$ as $\mathbb{C}^{\infty}_{\mathbb{C}}(B)$–bimodules. Then the maps
$f^{i}_{j}:P$ $\longrightarrow V,$
$x$ $\longrightarrow v=\sqrt{\rho_{b_{i}}(\pi(x))}\;\alpha(A(x)^{-1},e_{j}),$
where $A(x)$ is the unique element of $G$ such that $\Psi_{b_{i}}(\pi(x),A(x))=x$, from a set of $\mathbb{C}^{\infty}_{\mathbb{C}}(B)$–bimodule generators of $C^{\infty}_{\mathbb{C}}(P,V)^{G}$ associated to $\{s^{i}_{j}\}^{r,n}_{i,j=1}$. We define the smooth functions
$f^{i}_{jk}:P$ $\longrightarrow\mathbb{C}$
$x$ $\longrightarrow\langle e_{k}|f^{i}_{j}(x)\rangle,$
where $\langle$-|$\rangle$ - | - ⟩ is the inner product that makes $\delta^{V}$ unitary (antilinear in the second coordinate). Now, let us consider the linear maps
$T^{i}_{j}:V\longrightarrow C^{\infty}_{\mathbb{C}}(P)$
given by $T^{i}_{j}(e_{k})=f^{i}_{jk}$. A direct calculation shows that for all $x$in $P$ and for all $A$in $G$
$\left(\sum^{r,n}_{l,k=1}f^{l\,\ast}_{ki}f^{l}_{kj}\right)(x)=\delta_{ij},\qquad(\Delta_{P}\circ T^{i}_{j})(x,A)=((T^{i}_{j}\otimes\mathrm{id}_{H})\circ\delta^{V})(x,A).$
Then proposition follows by taking $T^{\mathrm{L}}_{1}=T^{1}_{1}$, $T^{\mathrm{L}}_{2}=T^{1}_{2}$,…, $T^{\mathrm{L}}_{n}=T^{1}_{n}$, $T^{\mathrm{L}}_{n+1}=T^{2}_{1}$,…, $T^{\mathrm{L}}_{d_{V}}=T^{r}_{n}$ with $d_{V}=rn$. ∎
It is worth mentioning that the maps $T^{i}_{j}$ agree with the dualization of $f^{i}_{j}$ via the pull–back, once the dual space $V^{\#}$ is identified with $V$. Moreover, to define the maps $\{T^{\mathrm{L}}_{i}\}^{d_{V}}_{i=1}$, it was necessary to have a partition of unity on the base space with smooth square roots. This is why having a $C^{\ast}$–algebra as the quantum base space (or one that can be completed to a $C^{\ast}$–algebra) is a sufficient condition to guarantee the existence of the maps $\{T^{\mathrm{L}}_{i}\}_{i=1}^{d_{V}}$ in the non–commutative geometrical setting, as the reader can verify in [8].
4.1. Quantum Principal Bundles and Dunkl Operators
For an example of the theory presented in [1], we will use somewhat lesser-known quantum bundles: the quantum principal $\mathcal{G}$–bundles developed in [6, 12]. We shall denote them by
$\zeta=(P,B,\Delta_{P}).$
These qpb’s are defined by the dualization of classical principal bundles with (finite) Coxeter groups $W(H)$ as the structure group, which we shall denote them by
$\rho:P_{\mathrm{class}}\longrightarrow P_{\mathrm{class}}/W(H),$
where $P_{\mathrm{class}}$ is the classical total space, $P_{\mathrm{class}}/W(H)$ is the classical base space, and the map $\rho:P_{\mathrm{class}}\longrightarrow P_{\mathrm{class}}/W(H)$ is the canonical projection (which is the bundle projection).
On the other hand, horizontal forms on $\zeta$ are given by the complexification of the de–Rham graded differential $\ast$–algebra of $P_{\mathrm{class}}$, quantum differential forms of $B$ are given by the complexification of the de–Rham graded differential $\ast$–algebra of $P_{\mathrm{class}}/W(H)$, and $(\Gamma,d)$ is given by the theory of $\ast$–FODC’s on finite groups [12]. It is worth mentioning that in $\zeta$ there exists a canonical qpc
(51) $\omega^{c}:\mathfrak{qg}^{\#}$ $\longrightarrow\Omega^{1}(P)$
$\theta$ $\longmapsto\mathbbm{1}_{B}\otimes\theta.$
and its covariant derivative $D^{\omega^{c}}$ is exactly the de-Rham differential. Since $\omega^{c}$ is real and regular, $\widehat{D}^{\omega^{c}}=D^{\omega^{c}}$.
In order to give a concrete example let us focus our study on standard Dunkl connections using the development in [12]. It is worth mentioning that, since we have changed the standard definition of qpc’s in order to embrace a more general theory, for us Dunkl displacements $\lambda:\mathfrak{qg}^{\#}\longrightarrow\Omega^{1}(P)$ do not need the $i=\sqrt{-1}$ factor. These qpc’s are given by
$\omega=\omega^{\mathrm{c}}+\lambda,$
where $\lambda$ satisfies $\lambda(\pi(\phi_{\sigma_{r}}))=\varrho_{r}r^{\#}$,
$\varrho_{r}:P_{\mathrm{class}}$ $\longrightarrow\mathbb{R}$
$x$ $\longrightarrow{\kappa(r)\over\langle r|x\rangle},$
$\kappa:R\longrightarrow\mathbb{R}$ is a multiplicative function (it is $W(H)$–invariant), $r$in $R$ with $R$ the corresponding root system, and $r^{\#}$ is the element of the dual space associated with $r$. Covariant derivatives of this kind of qpc’s are given by
$(D^{\omega}f)(x)=df(x)+\sum_{r\in R^{+}}\kappa(r){f(x)-f(x\sigma_{r})\over\langle r|x\rangle}r$
for every $f$in $P$, which is a Dunkl operator in vector form [6].
Let $\delta^{V}$in $\mathcal{T}$. Then
$\nabla^{\omega}_{V}T=\sum^{n_{V}}_{i=1}\mu^{D^{\omega}T}_{i}\otimes_{B}T^{V}_{i}\,,\quad\widehat{\nabla}^{\omega}_{V}T=\sum^{n_{V}}_{i=1}T^{V}_{i}\otimes_{B}\mu^{D^{\omega}T}_{i},$
where $\mu^{D^{\omega}T}_{k}=\displaystyle\sum^{n_{V}}_{i=1}D^{\omega}T(e_{i})f^{V\,\ast}_{ki}$ and $T(e_{k})$in $P$ [1]. Notice that Dunkl connections are multiplicative, but not regular [6]. Furthermore, it is worth mentioning that, in differential geometry, there are no principal connections on the principal bundle $\rho:P_{\mathrm{class}}\longrightarrow P_{\mathrm{class}}/W(H)$.
On the other hand, the left and right canonical Hermitian structures are given by
$\langle T_{1},T_{2}\rangle_{\mathrm{L}}=\sum^{n_{V}}_{k=1}T_{1}(e_{k})T_{2}(e_{k})^{\ast}\,,\qquad\langle T_{1},T_{2}\rangle_{\mathrm{R}}=\sum^{n_{V}}_{k=1}T_{1}(e_{k})^{\ast}T_{2}(e_{k}).$
For an explicit example of Theorem $3.14$ of [1], let us take a real Dunkl connection. These connections are characterized by $\omega=\omega^{\mathrm{c}}+\widetilde{\lambda}$, with $\widetilde{\lambda}=i\lambda$. Therefore, using equation (30),
$\langle\nabla^{\omega}_{V}(T_{1}),T_{2}\rangle_{\mathrm{L}}\,=\,\sum^{n_{V}}_{i=1}\mu^{D^{\omega}T_{1}}_{i}\langle T^{V}_{i},T_{2}\rangle_{\mathrm{L}}\,=\,\sum^{n_{V}}_{k=1}\left(D^{\omega}T_{1}(e_{k})\right)T_{2}(e_{k})^{\ast};$
and evaluating on any $x$in $P_{\mathrm{class}}$ we have
$\sum^{n_{V}}_{k=1}\left(dT_{1}(e_{k})(x)T_{2}(e_{k})^{\ast}(x)+i\sum_{r\in R^{+}}\kappa(r){T_{1}(e_{k})(x)-T_{1}(e_{k})(x\sigma_{r})\over\langle r|x\rangle}r\,T_{2}(e_{k})^{\ast}(x)\right).$
In the same way
$\langle T_{1},\nabla^{\omega}_{V}(T_{2})\rangle_{\mathrm{L}}\,=\,\sum^{n_{V}}_{i=1}\langle T_{1},T_{i}\rangle_{\mathrm{L}}\left(\mu^{D^{\omega}T_{2}}_{i}\right)^{\ast}\,=\,\sum^{n_{V}}_{k=1}T_{1}(e_{k})\left(D^{\omega}T_{2}(e_{k})\right)^{\ast};$
and evaluating on any $x$in $P_{\mathrm{class}}$ we get
$\sum^{n_{V}}_{k=1}\left(T_{1}(e_{k})(x)dT_{2}(e_{k})^{\ast}(x)-i\,T_{1}(e_{k})(x)\sum_{r\in R^{+}}\kappa(r){T_{2}(e_{k})^{\ast}(x)-T_{2}(e_{k})^{\ast}(x\sigma_{r})\over\langle r|x\rangle}r\right).$
This implies for all $x$in $P_{\mathrm{class}}$ that $\langle\nabla^{\omega}_{V}(T_{1}),T_{2}\rangle_{\mathrm{L}}+\langle T_{1},\nabla^{\omega}_{V}(T_{2})\rangle_{\mathrm{L}}$ is equal to
$d\langle T_{1},T_{2}\rangle_{\mathrm{L}}(x)+i\sum_{r\in R^{+}}\kappa(r){q(x)-q(x\sigma_{r})\over\langle r|x\rangle}r,$
where $q=\displaystyle\sum^{n_{V}}_{k=1}T_{1}(e_{k})(\sigma_{r}\cdot T_{2}(e_{k}))^{\ast}$in $P$. However, since $T_{1}$, $T_{2}$in $\textsc{Mor}(\delta^{V},\Delta_{P})$, we have $\sigma_{r}\cdot T_{i}(e_{j})=\displaystyle\sum^{n_{V}}_{k=1}\lambda^{\sigma_{r}}_{kj}T_{i}(e_{k})$. Thus $q=\displaystyle\sum^{n_{V}}_{k=1}(\sigma_{r}\cdot T_{1}(e_{k}))T_{2}(e_{k})^{\ast}$ and hence $q(x)=q(x\sigma_{r})$ for all $x$in $P_{\mathrm{class}}$. In summary, we have proven explicitly that
$\langle\nabla^{\omega}_{V}(T_{1}),T_{2}\rangle_{\mathrm{L}}+\langle T_{1},\nabla^{\omega}_{V}(T_{2})\rangle_{\mathrm{L}}=d\langle T_{1},T_{2}\rangle_{\mathrm{L}}.$
A similar argument proves that
^$\langle\widehat{\nabla}^{\omega}_{V}(T_{1}),T_{2}\rangle_{\mathrm{R}}+\langle T_{1},\widehat{\nabla}^{\omega}_{V}(T_{2})\rangle_{\mathrm{R}}=d\langle T_{1},T_{2}\rangle_{\mathrm{R}}.$
Defining $g^{V}_{ij}:W(H)\longrightarrow\mathbb{C}$ by $g^{V}_{ij}(h)=\lambda^{h}_{ij}$ we get
$\mathrm{qtrs}(g^{V}_{ij})=\sum^{n_{V}}_{k=1}f^{V\,\ast}_{ki}\otimes_{B}f^{V}_{kj},$
and by considering $\omega^{c}$ in equation (37), we have for all $\theta$in $\mathfrak{qg}^{\#}$
(52) $\mathrm{qtrs}(\theta)=\mathbbm{1}\otimes_{\Omega^{\bullet}(B)}(\mathbbm{1}\otimes\theta)-(\mathbbm{1}\otimes\theta^{(0)})\,\mathrm{qtrs}(\theta^{(1)}),$
where $\mathrm{ad}(\theta)=\theta^{(0)}\otimes\theta^{(1)}$.
Now let us consider a graded left $\Omega^{\bullet}(B)$–module isomorphism
$\mathfrak{F}:\Omega^{\bullet}(B)\longrightarrow\Omega^{\bullet}(B)$
which (for irreducible degree $1$ elements) is given by
$\mathfrak{F}(\mu+x\otimes\theta)=\mu+x\lambda(\theta)+x\otimes\theta,$
where $\lambda$ is a Dunkl displacement, $\mu$in $\mathrm{Hor}^{1}P$, $x$in $P$, and $\theta$in $\mathfrak{qg}^{\#}$. This map is actually a qgt $\mathfrak{f}$ and a direct calculation shows that $\mathfrak{F}^{\circledast}\omega^{c}$ is a standard Dunkl connection. By using the form of the last map $\mathfrak{F}$, every single classical gauge transformation can be extended into a quantum one.